### v1

# A Nonabelian Hodge Correspondence for Principal Bundles in Positive Characteristic

Abstract.
In this paper, we prove a nonabelian Hodge correspondence for principal bundles on a smooth variety $X$ in positive characteristic, which generalizes the Ogus-Vologodsky correspondence for vector bundles. Then we extend the correspondence to logahoric torsors over a log pair $(X,D)$, where $D$ a reduced normal crossing divisor in $X$. As an intermediate step, we prove a correspondence between principal bundles on root stacks $\mathscr{X}$ and parahoric torsors on $(X,D)$, which generalizes the correspondence on curves given by Balaji–Seshadri to higher dimensional case.
11footnotetext: Key words: nonabelian Hodge correspondence, G-Higgs bundle, flat G-bundle22footnotetext: MSC2020: 14C30, 14L15, 20G15
1. Introduction
Ogus–Vologodsky established the nonabelian Hodge correspondence for $G={\rm GL}_{n}$ in positive characteristic [OV07]. A natural question is how to generalize Ogus–Vologodsky’s result to an arbitrary reductive group $G$. On the other hand, Chen–Zhu established a nonabelian Hodge correspondence (not in one to one manner) for reductive groups on curves with the help of Hitchin morphisms [CZ15]. For nilpotent objects with small exponent their correspondence coincides with the Ogus-Vologodsky correspondence. However, their approach seems to be hard to be generalized to higher dimensional varieties, because there is still a lack of numerous necessary properties of the Hitchin morphism (see [CN20, Conjecture 3.2 and Conjecture 5.2]).
As remarked in [LSZ15, Remark 2.4], the exponential twisting approach to the Ogus-Vologodsky correspondence should respect $G$-structure for any subgroup $G\subset\mathrm{GL}$. In this paper, we aim to work out this point in detail, which is going to be applied to construct many absolutely irreducible crystalline $G$-local systems in a subsequent work. As would be clear to the reader, the key point in this approach is the construction of exponential maps for reductive groups with good properties. Fortunately, this has been studied by Serre [Ser94] and several other authors in a series work [Sei00, McN02, Sob15]. In particular, Sobaje gave a definition of the exponential map in positive characteristic [Sob15, Definition 2.3] and prove its existence and uniqueness when the characteristic is pretty good [Sob15, Theorem 3.4]. These knowledge suffice for generalizing the result of Lan–Sheng–Zuo [LSZ15] in the vector bundle case to principal $G$-bundles, where $G$ denotes for a connected split reductive group $G$ over a perfect field $k$ in char $p>0$. In effect, we establish a correspondence between nilpotent $G$-Higgs bundles and nilpotent flat $G$-bundles of exponent $\leq p-1$ for a smooth variety over $k$ which is $W_{2}(k)$-liftable. Next, we generalize this correspondence to the special log setting $(X,D)$, where $X$ is a smooth variety and $D$ is a reduced normal crossing divisor of $X$. There has been the work by Schepler [Sch05] and the work [LSYZ19] via exponential twisting in the vector bundle case, and the work by Krishnamoorthy and the first named author [KS20] for parabolic bundles over curves. The right generalization to principal bundles of parabolic Higgs bundles which have simple poles along $D$ is the notion of logahoric (=logarithmic parahoric) Higgs torsors, which was proposed by P. Boalch [Boa11] in characteristic zero (see also [KSZ24]). Like in [KS20], we first establish a logarithmic version of the correspondence over a tame root stack. However, the procedure of descending logarithmic Higgs torsors/connections from root stacks to logahoric Higgs torsors/connections over $(X,D)$ is bit different from the generalized Biswas-Iyer-Simpson correspondence for parabolic vector bundles in loc. cit. When $X$ is projective equipped with an ample line bundle $H$, one has the notion of R-stability [Ram75, Ram96a] for principal bundles which generalizes the Mumford-Takemoto slope stability in vector bundle case [Ram75, §3]. It has been further generalized to a log pair $(X,D)$ in [KSZ24]. We show that our correspondence preserves the R-stability condition, as one may expect.
Now we give the setup and main results in this paper. Let $k$ be a perfect field in positive characteristic $p$. Let $X$ be a $W_{2}(k)$-liftable smooth variety over $k$ and denote by $X^{\prime}$ the Frobenius pullback. Let $G$ be a connected split reductive linear algebraic group over $k$ and we suppose that the characteristic $p$ is pretty good [Her13, Definition 2.11]. We introduce the following categories
  • ${\rm HIG}_{p-1}(X^{\prime}/k,G)$: the category of nilpotent $G$-Higgs bundles of exponent $\leq p-1$ on $X$;
  • ${\rm MIC}_{p-1}(X/k,G)$: the category of nilpotent flat $G$-bundles of exponent $\leq p-1$ on $X$,
and prove that they are equivalent by establishing the exponential twisting of the Cartier and inverse Cartier transform in §3.
Theorem 1.1 (Theorem 3.1).
Suppose that $X$ is $W_{2}(k)$-liftable. The categories ${\rm HIG}_{p-1}(X^{\prime}/k,G)$ and ${\rm MIC}_{p-1}(X/k,G)$ are equivalent.
Suppose $X$ is projective and we fix an ample line bundle $H$ on $X$. We follow Ramanathan’s approach [Ram75, Ram96a] to define the stability condition of $G$-bundles on $X$ with an arbitrary dimension, which is called $R$-stability condition (see §2.3). Denote by ${\rm HIG}^{s(s)}_{p-1}(X^{\prime}/k,G)\subseteq{\rm HIG}_{p-1}(X^{\prime}/k)$ (resp. ${\rm MIC}^{s(s)}_{p-1}(X/k,G)\subseteq{\rm MIC}_{p-1}(X/k)$) the full subcategory of $R$-(semi)stable nilpotent $G$-Higgs bundles (resp. flat $G$-bundles) of exponent $\leq p-1$. We prove that the correspondence also holds under $R$-stability conditions.
Theorem 1.2 (Theorem 3.6).
The categories ${\rm HIG}^{s(s)}_{p-1}(X^{\prime}/k,G)$ and ${\rm MIC}^{s(s)}_{p-1}(X/k,G)$ are equivalent.
Now we consider the correspondence on $X$ together with a reduced normal crossing divisor $D=\sum_{i=1}^{s}D_{i}$, where $D_{i}$ are irreducible components for $1\leq i\leq s$. Following Simpson’s idea [Sim90], the analogue of principal bundles on $X\backslash D$ is parahoric torsors [HKSZ22], especially when discussing stability conditions. In a nutshell, we prove the equivalence of categories between nilpotent logahoric Higgs torsors and nilpotent logahoric connections of exponent $\leq p-1$. Here is a sketch of the proof. We equip each $D_{i}$ with a tame weight $\alpha_{i}$, where a tame weight is a cocharacter with rational coefficients such that the denominator $d_{i}$ is coprime to $p$. Denote by $\boldsymbol{\alpha}$ (resp. $\boldsymbol{d}$) the collection of weights (resp. integers). With respect to the above data, people constructed a (tame) root stack $\mathscr{X}$ ([MO05, Theorem 4.1]) and a smooth affine group scheme $\mathcal{G}_{\boldsymbol{\alpha}}$ on $X$ [BP22, Theorem 5.4]. We define the following categories
  • ${\rm HIG}_{p-1}(\mathscr{X}_{\rm log}/k,G)$: the category of nilpotent logarithmic $G$-Higgs bundles of exponent $\leq p-1$ on $\mathscr{X}$;
  • ${\rm MIC}_{p-1}(\mathscr{X}_{\rm log}/k,G)$: the category of nilpotent logarithmic flat $G$-bundles of exponent $\leq p-1$ on $\mathscr{X}$, of which the residues are also nilpotent of exponent $\leq p-1$.
We first prove a logarithmic version of the nonabelian Hodge correspondence on the root stack $\mathscr{X}$ by applying results in §3 and [LSYZ19, Appendix].
Theorem 1.3 (Theorem 4.12).
Suppose that $(\mathscr{X},\widetilde{D})$ is $W_{2}$-liftable. The categories ${\rm HIG}_{p-1}(\mathscr{X}^{\prime}_{\rm log}/k,G)$ and ${\rm MIC}_{p-1}(\mathscr{X}_{\rm log}/k,G)$ are equivalent.
Similar to the case of curves [BS15, §3], we prove that parahoric $\mathcal{G}_{\boldsymbol{\alpha}}$-torsors on $X$ are in one-to-one correpsondence with $G$-bundles on $\mathscr{X}$ of type $\boldsymbol{\rho}$, where $\boldsymbol{\rho}=\{\rho_{i},1\leq i\leq s\}$ is a collection of representations $\rho_{i}:\mu_{d_{i}}\rightarrow T$ determined by $\boldsymbol{\alpha}$ and $T$ is a maximal torus of $G$.
Proposition 1.4 (Proposition 5.5).
The category of parahoric $\mathcal{G}_{\boldsymbol{\alpha}}$-torsors on $X$ and the category of $G$-bundles on $\mathscr{X}$ of type $\boldsymbol{\rho}$ are equivalent.
This correspondence carries over $G$-bundles with Higgs field/connection. It has been done in [KSZ24] for curves in characteristic zero, and in [KS20] for vector bundles over curves in positive characteristic. We first introduce some definitions. A logahoric $\mathcal{G}_{\boldsymbol{\alpha}}$-Higgs torsor is a pair $(\mathcal{E},\vartheta)$, where $\mathcal{E}$ is a parahoric $\mathcal{G}_{\boldsymbol{\alpha}}$-torsor and $\vartheta:X\rightarrow\mathcal{E}(\mathfrak{g})\otimes\Omega_{X}({\rm log}\,D)$ is a logarithmic Higgs field. A logahoric $\mathcal{G}_{\boldsymbol{\alpha}}$-connection is a pair $(\mathcal{V},\nabla^{\prime})$, where $\mathcal{V}$ is a parahoric $\mathcal{G}_{\boldsymbol{\alpha}}$-torsor and $\nabla^{\prime}:\mathcal{O}_{\mathcal{V}}\rightarrow\mathcal{O}_{\mathcal{V}}\otimes\Omega_{X}({\rm log}\,D)$ is a logarithmic integrable connection. We define
  • ${\rm HIG}_{p-1}(X_{\rm log}/k,\mathcal{G}_{\boldsymbol{\alpha}})$: the category of nilpotent logarhoic $\mathcal{G}_{\boldsymbol{\alpha}}$-Higgs torsors of exponent $\leq p-1$ on $X$;
  • ${\rm MIC}_{p-1}(X_{\rm log}/k,\mathcal{G}_{\boldsymbol{\alpha}})$: the category of nilpotent logahoric $\mathcal{G}_{\boldsymbol{\alpha}}$-connections of exponent $\leq p-1$ on $X$ such that the semisimple parts of residues are $\boldsymbol{\alpha}$.
Note that, in the above definition, the residue of a logarithmic connection on a parahoric torsor is not nilpotent and the semisimple part is given by the weights $\boldsymbol{\alpha}$.
Proposition 1.5 (Corollary 5.12 and Proposition 5.14).
The categories
  • ${\rm HIG}_{p-1}(X_{\rm log}/k,\mathcal{G}_{\boldsymbol{\alpha}})$ and ${\rm HIG}_{p-1}(\mathscr{X}_{\rm log}/k,G,\boldsymbol{\rho})$,
  • ${\rm MIC}_{p-1}(X_{\rm log}/k,\mathcal{G}_{\boldsymbol{\alpha}})$ and ${\rm MIC}_{p-1}(\mathscr{X}_{\rm log}/k,G,\boldsymbol{\rho})$
are equivalent.
Furthermore, the above equivalences also hold under $R$-stability conditions for Higgs bundles and flat $G$-bundles. Therefore, we follow the diagram below and prove a logahoric version of nonabelian Hodge correspondence:
[…]
where $\boldsymbol{\rho}^{p}=\{\rho_{i}^{p},1\leq i\leq s\}$ and $\boldsymbol{\alpha}^{\prime}$ is a collection of weights determined by $\boldsymbol{\rho}^{p}$.
Theorem 1.6 (Theorem 5.17).
Suppose that $(X,D)$ is $W_{2}(k)$-liftable. The categories ${\rm HIG}_{p-1}(X^{\prime}_{\rm log}/k,\mathcal{G}_{\boldsymbol{\alpha}})$ and ${\rm MIC}_{p-1}(X_{\rm log}/k,\mathcal{G}_{\boldsymbol{\alpha}^{\prime}})$. Moreover, the equivalence preserves stability conditions.
We make some further remarks and explain relations to previous work as follows.
Ogus–Vologodsky established the nonabelian Hodge correspondence in positive characteristic for $G={\rm GL}_{n}$ [OV07] and the logarithmic case was studied by Schepler [Sch05]. Our work is a generalization of Ogus–Vologodsky’s correspondence for general reductive groups and we also extend the results to parahoric torsors. Also, Chen–Zhu introduced an analogue of Hitchin’s equations for principal bundles in positive characteristic
$\begin{cases}\psi(\nabla-\theta)=0\\ ((\nabla-\theta)\otimes\nabla^{can})(\psi(\nabla))=0\end{cases}.$
where $\nabla$ is a connection, $\theta$ is a Higgs field and $\psi$ is the $p$-curvature [CZ15, Introduction]. The main results in this paper (Theorem 3.1 and Theorem 5.17) provide solutions for these equations in higher dimensional case.
Parahoric torsors were studied by many groups of people. Heinloth studied uniformization property of parahoric torsors [Hei10]. After that, Balaji–Seshadri constructed the moduli space of parahorc torsors in the case of curves [BS15]. Recently, Balaji–Pandey defined parahoric group schemes over higher dimensional bases [BP22]. Similar to the case of curves [BS15, §3], parahoric torsors are equivalent to principal bundles on an appropriate root stacks. This property is observed by Balaji–Pandey locally [BP22] and is proven in §5.2.
The $R$-stability condition for $G$-bundles is introduced by Ramanathan in the case of curves [Ram75, Ram96a]. When $G$ is a classical group, the $R$-stability condition of a principal bundle is equivalent to the slope stability condition of the corresponding associated bundle in the case of curves. We refer the reader to [CS21, §6] and [Yan22] for more details. In a similar way, Heinloth define a stability condition for parahoric torsors [Hei17, §1]. The stability condition he considered is equivalent to the slope stability condition, which does not include the information of weights. On the other hand, the stability condition considered for parahoric torsors in this paper include the information of weights (Definition 5.6 and 5.7), which is equivalent to the stability condition of the corresponding parabolic bundles when $G={\rm GL}_{n}$ [KSZ24, §5]. Moreover, it is also equivalent to the one given by Balaji-Seshadri [BS15]. Moreover, as proved in this paper, the stability condition of parahoric torsors is equivalent to the stability condition of the corresponding $G$-bundles on root stacks (Proposition 5.9).
In the parahoric case, the residues of nilpotent logarithmic connections are not nilpotent in general and the semisimple part of the residue is given by the corresponding weight. This phenomenon actually comes from the establishment of the equivalence between parahoric torsors on $X$ and principal bundles on the root stack $\mathscr{X}$. We introduce a new element $\Delta_{i}$ for each $1\leq i\leq s$ and define new transition functions. The change of transition functions results in the change of residues. In positive characteristic, this phenomenon first appears in [KS20], where the authors studied the case of $G={\rm GL}_{n}$ and named the condition as adjustedness [KS20, Definition 2.9]. Later on, the authors generalized the result to parahoric torsors on curves [LS24], which also generalized Chen–Zhu’s result [CZ15]. Moreover, in characteristic zero, a similar discussion is also given in [Sim90, §5].
2. Preliminaries
In this section, we introduce the main objects considered in this paper: (semistable) nilpotent $G$-Higgs bundles and (semistable) nilpotent flat $G$-bundles. In §2.1, we review definitions of $G$-Higgs bundles and flat $G$-bundles, and a very good reference is [CZ15, Appendix]. In §2.2, we introduce the nilpotency and review the exponential map for linear algebraic groups given in [Sob15]. In §2.3, we give the definition of stability conditions, which come from Ramanathan’s definition for principal bundles [Ram75, Ram96a] and are called $R$-stability conditions. We also refer the reader to [KSZ24] for more details about the case in characteristic zero.
2.1. G-Higgs bundles and Flat G-bundles
Let $X$ be a smooth variety over $k$, which is a perfect field in positive characteristic $p$. Denote by $T_{X/k}$ (resp. $\Omega_{X/k}$) the relative tangent (resp. cotangent) sheaf. Sometimes we use the notation $T_{X}$ and $\Omega_{X}$ if there is no ambiguity. Since we are working in characteristic $p$, we have the following well-known diagram
[…]
where
  • $X^{\prime}:=X\times_{{\rm Spec}\,k,F}{\rm Spec}\,k$,
  • $F_{k}$ (resp. $F_{X}$) is the absolute Frobenius morphism of ${\rm Spec}\,k$ (resp. $X$),
  • $F_{X/k}$ is the relative Frobenius morphism.
Let $G$ be a connected split reductive linear algebraic group over $k$ with maximal torus $T$. Denote by $\mathcal{R}$ the set of roots. Let $\mathfrak{g}$ (resp. $\mathfrak{t}$) be the Lie algebra of $G$ (resp. $T$). Let $E$ be a $G$-bundle on $X$ with adjoint bundle $E(\mathfrak{g}):=E\times_{G}g$. The Lie bracket [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] of $\mathfrak{g}$ induces a well-defined binary operation
$E(\mathfrak{g})\otimes_{\mathcal{O}_{X}}E(\mathfrak{g})\rightarrow E(\mathfrak{g}),$
and we use the same notation [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] for this morphism (see [BBP17, §8] for instance). Furthermore, if the tensor product is defined over $\mathcal{O}_{X}$, we use the notation $E(\mathfrak{g})\otimes E(\mathfrak{g})$ for convenience. Now given a section $\theta\in H^{0}(X,E(\mathfrak{g})\otimes\Omega_{X})$, the composition of the following maps
$X\xrightarrow{\theta\otimes\theta}(E(\mathfrak{g})\otimes\Omega_{X})\otimes(E(\mathfrak{g})\otimes\Omega_{X})\rightarrow(E(\mathfrak{g})\otimes E(\mathfrak{g})\otimes(\Omega_{X}\otimes\Omega_{X})\xrightarrow{[\cdot,\cdot]\otimes\wedge}\rightarrow E(\mathfrak{g})\otimes\Omega^{2}_{X}$
is denoted by $\theta\wedge\theta$. We say that $\theta$ is integrable if $\theta\wedge\theta=0$, and an integrable section is called a Higgs field.
Definition 2.1.
A $G$-Higgs bundle on $X$ is a pair $(E,\theta)$, where $E$ is a $G$-bundle on $X$ and $\theta$ is a Higgs field.
Now we consider connections on $G$-bundles. There is a canonical connection on $\mathcal{O}_{G\times X}$ and denote it by $\nabla_{G}$. Let $V$ be a $G$-bundle. A (integrable) $G$-connection $\nabla$ on $V$ is a (integrable) connection $\nabla:\mathcal{O}_{V}\rightarrow\mathcal{O}_{V}\otimes_{\mathcal{O}_{X}}\Omega_{X}$ such that the following diagram commutes
[…]
where $a$ is the co-action map. Here is an equivalent definition by Lie algebroids. Let ${\rm At}(V)$ be the Atiyah algebroid of $V$ (see [Bis10] or [CZ15, Appendix]). There is a natural projection ${\rm At}(V)\rightarrow T_{X}$ together with a short exact sequence
$0\rightarrow V(\mathfrak{g})\rightarrow{\rm At}(V)\rightarrow T_{X}\rightarrow 0$
A $G$-connection $\nabla:T_{X}\rightarrow{\rm At}(V)$ is a splitting of the exact sequence, and it is integrable if the splitting is a homomorphism of Lie algebroids.
Definition 2.2.
A flat $G$-bundle is a pair $(V,\nabla)$, where $V$ is a $G$-bundle and $\nabla$ is an integrable $G$-connection.
Example 2.3.
This example is discussed in [CZ15, Appendix]. Since it is very helpful to understand Higgs fields and connections locally, we also include it here. Let $V$ be a trivial $G$-bundle on $X$, and the Atiyah algebroid ${\rm At}(V)\cong V(\mathfrak{g})\oplus T_{X}$ is a direct sum. Then a connection $\nabla$ is equivalent to a morphism $T_{X}\rightarrow V(\mathfrak{g})$, and therefore, the sheaf of integrable connections on $X$ is isomorphic to a subsheaf of $V(\mathfrak{g})\otimes\Omega_{X}$ and denote it by $(V(\mathfrak{g})\otimes\Omega_{X})^{cl}$, where the upperscript is for the word “closed”. The corresponding element of a given connection in $(V(\mathfrak{g})\otimes\Omega_{X})^{cl}$ is usually regarded as the connection form.
Now let $V$ be an arbitrary $G$-bundle. We can find a covering $\mathcal{U}=\{U_{i}\}_{i\in I}$, where $I$ is the index set, such that $V|_{U_{i}}\cong G\times U_{i}$. Therefore, an integrable $G$-connection $\nabla$ on $V$ can be regarded as a section of $(V|_{U_{i}}(\mathfrak{g})\otimes\Omega_{U_{i}})^{cl}$ locally. This approach shows that given a Higgs field $\theta$ and an integrable $G$-connection $\nabla$, the sum $\nabla+\theta$ is a well-defined integrable $G$-connection locally.
2.2. Nilpotency
In this subsection, we define nilpotent $G$-Higgs bundles and nilpotent flat $G$-bundles. We first make a brief review of distributions and exponentiations. We refer the reader to [Jan87, I.7] for more details about distributions and [Ser94, Sei00] about exponentiations.
Distribution
Let $e\in G$ be the identity element. Denote by $I\subseteq k[G]$ the ideal $I=\{f\in k[G]\text{ }|\text{ }f(e)=0\}$. A distribution on $G$ with support in $e$ of order $\leq n$ is a $k$-linear map $k[G]\rightarrow k$, which maps $I^{n+1}$ to zero. Denote by ${\rm Dist}_{n}(G)$ the set of all distributions on $G$ with support in $e$ of order $\leq n$. Clearly, we have
${\rm Dist}_{n}(G):=(k[G]/I^{n+1})^{*}.$
Let ${\rm Dist}^{+}_{n}(G)\subseteq{\rm Dist}_{n}(G)$ consists of all distributions which send $e$ to zero, and then
${\rm Dist}^{+}_{n}(G)=(I/I^{n+1})^{*},\,{\rm Dist}_{n}(G)\cong k\oplus{\rm Dist}^{+}_{n}(G).$
There is a natural way to define a multiplication on distributions such that
${\rm Dist}_{n}(G)\cdot{\rm Dist}_{m}(G)\subseteq{\rm Dist}_{n+m}(G).$
Then ${\rm Dist}(G):=\bigcup_{n\geq 0}{\rm Dist}_{n}(G)$ is a $k$-algebra, which is called the algebra of distributions on $G$. Moreover, given $f\in{\rm Dist}_{n}(G)$ and $g\in{\rm Dist}_{m}(G)$, we have
$[f,g]=fg-gf\in{\rm Dist}_{n+m-1}(G).$
In particular, if $f,g\in{\rm Dist}^{+}_{1}(G)$, we get
$[f,g]=fg-gf\in{\rm Dist}^{+}_{1}(G),$
which gives a Lie algebra structure on ${\rm Dist}^{+}_{1}(G)$, and ${\rm Dist}^{+}_{1}(G)\cong\mathfrak{g}$. Therefore, given elements $a,b\in\mathfrak{g}$, the multiplication $ab$ is defined as the multiplication of the corresponding distributions. As a special case, the $p$-th power of an arbitrary element $a\in\mathfrak{g}\cong{\rm Dist}^{+}_{1}(G)$ also lies in $\mathfrak{g}$, and we denote it by $a^{[p]}$. If $a^{[p]}=0$, $a$ is called $[p]$-nilpotent. Moreover, if $a$ is $[p]$-nilpotent, then $a$ is nilpotent of exponent $\leq p-1$ because $({\rm ad}(a))^{p}=0$ in characteristic $p$, where ${\rm ad}(a):\mathfrak{g}\rightarrow\mathfrak{g}$ is the adjoint operator.
Exponentiation
Given a connected reductive split reductive linear algebraic group $G$, we suppose that the characteristic $p$ is pretty good. We refer the reader to [Her13, Definition 2.11] for the definition and [Sob15, Introduction] for examples. Roughly speaking, if $G$ is not of type $A$, $p$ is pretty good if $p>7$, while if $G$ is of type $A_{n}$, $p$ should satisfy the conditions that $p>2$ and $p$ is not divided of $n$. Sobaje studied the exponentiation of commuting nilpotent varieties and introduced the following definition:
Definition 2.4 (Definition 2.3 in [Sob15]).
Let $H$ be an affine algebraic group over $k$. We say that $H$ has an exponential map if there exists an $H$-equivariant injective map $\overline{\rm exp}:\mathcal{N}_{1}(\mathfrak{h})\rightarrow H$, where $\mathcal{N}_{1}(\mathfrak{h})$ is the subvariety of $[p]$-nilpotent element in the Lie algebra $\mathfrak{h}$, such that for each $a\in\mathcal{N}_{1}(\mathfrak{h})$, we have
  1. (1)
    the map $\overline{\rm exp}$, which sends $t\in\mathbb{G}_{a}$ to $\overline{\rm exp}(ta)$, is a one-parameter additive subgroup of $H$ whose differential sends $\frac{\partial}{\partial t}$ to $a$;
  2. (2)
    the one-parameter subgroup $\overline{\rm exp}(\mathbb{G}_{a})$ acts trivially (as adjoint action) on $C_{\mathfrak{h}}(a):=\{b\,|\,[a,b]=0\}\subseteq\mathfrak{h}$, i.e.
    ${\rm Ad}(\overline{\rm exp}(a))(b)=b$
    for $b\in C_{\mathfrak{h}}(a)$.
Based on this definition, Sobaje proved the following theorem.
Theorem 2.5 (Theorem 3.4 in [Sob15]).
Let $\mathcal{N}_{1}(\mathfrak{g})$ be the $[p]$-nilpotent variety of $\mathfrak{g}$ and let $\mathcal{U}_{1}(G)$ be the $p$-unipotent variety of $G$. Then there is a unique $G$-equivariant bijection
$\overline{\rm exp}:\mathcal{N}_{1}(\mathfrak{g})\xrightarrow{\cong}\mathcal{U}_{1}(G),$
which is an exponential map, satisfying the property that if $U$ is the unipotent radical of a parabolic subgroup $P\subseteq G$ such that $U$ has nilpotency class less than $p$, then $\overline{\rm exp}$ restricts to a $P$-equivariant isomorphism $\mathfrak{u}\xrightarrow{\cong}U$ of algebraic groups, where $\mathfrak{u}$ is the Lie algebra of $U$, having tangent maps equal to the identity on $\mathfrak{u}$.
Indeed, when the nilpotency of $U$ is less than $p$, the $P$-equivariant isomorphism $\mathfrak{u}\xrightarrow{\cong}U$ is first given by Serre [Ser94, §4] and then studied in [Sei00, Proposition 5.2] and [McN02, §8]. The exponential map given by Sobaje is crucial when we studied Cartier and inverse Cartier transform for nilpotent $G$-Higgs bundles and nilpotent flat $G$-bundles in this paper.
Nilpotent G-Higgs bundles and nilpotent flat G-bundles
Let $U$ be an open affine subset of $X$ such that $E|_{U}$ is a trivial $G$-bundle. Taking a section $\partial$ of $T_{U}$, $\theta|_{U}(\partial)$ is a section of $E|_{U}(\mathfrak{g})$. Since $E|_{U}(\mathfrak{g})\cong U\times\mathfrak{g}$, $\theta|_{U}(\partial)$ is regarded as an element in $\mathfrak{g}(\mathcal{O}_{U})$. A $G$-Higgs bundle $(E,\theta)$ is nilpotent of exponent $\leq p-1$, if there exists a covering of $X$ by open affine subsets $U$ such that for all local sections $\partial_{1},\dots,\partial_{p}$ of $T_{U}$, we have
$\theta|_{U}(\partial_{1})\cdots\theta|_{U}(\partial_{p})=0,$
where the product is defined as the multiplication of distributions. Let ${\rm HIG}_{p-1}(X/k,G)$ be the category of nilpotent $G$-Higgs bundles on $X$ of exponent $\leq p-1$. Furthermore, if $G={\rm GL}_{n}(k)$, we use the notation ${\rm HIG}_{p-1}(X/k)$ for the category of nilpotent Higgs bundles on $X$ of exponent $\leq p-1$.
Let $(V,\nabla)$ be a flat $G$-bundle on $X$. For any element $v\in T_{X}$, we define a map of sheaves $T_{X}\rightarrow\mathcal{E}nd(\mathcal{O}_{V})$ as $v\rightarrow\nabla(v)^{p}-\nabla(v^{[p]})$, where $v^{[p]}$ is $p$-th power of $v$. By [Kat70, §5], we obtain a morphism $\psi(\nabla):F^{*}_{X/k}T_{X^{\prime}}\rightarrow\mathcal{E}nd(\mathcal{O}_{V})$ of $\mathcal{O}_{X}$-modules, and for any $v\in F^{*}_{X/k}T_{X^{\prime}}$, the following diagram commutes by ($\ast$2.1).
${\mathcal{O}_{V}}$${\mathcal{O}_{V}\otimes_{\mathcal{O}_{X}}\mathcal{O}_{G\times X}}$${\mathcal{O}_{V}}$${\mathcal{O}_{V}\otimes_{\mathcal{O}_{X}}\mathcal{O}_{G\times X}}$$\scriptstyle{a}$$\scriptstyle{\psi(\nabla)(v)}$$\scriptstyle{\psi(\nabla)(v)\otimes 1+1\otimes\psi(\nabla_{G})(v)}$$\scriptstyle{a\otimes 1}$
This morphism $\psi(\nabla)$ is called the $p$-curvature of $(V,\nabla)$. If there is no ambiguity, we use the notation $\psi$ for the $p$-curvature. Since $\psi(\nabla_{G})=0$, the $p$-curvature $\psi(\nabla)$ can be regarded as a section of $V(\mathfrak{g})\otimes F^{*}_{X/k}\Omega_{X^{\prime}}$, which is called a $F$-Higgs field (see [CZ15, A.6]).
Given a flat $G$-bundle $(V,\nabla)$, it is nilpotent of exponent $\leq p-1$ if there exists a covering of $X$ by open affine subsets $U$ such that for all local sections $\partial_{1},\dots,\partial_{p}$ of $F^{*}_{U^{\prime}/k}T_{U^{\prime}}$, we have
$\psi|_{U}(\partial_{1})\cdots\psi|_{U}(\partial_{p})=0.$
Let ${\rm MIC}_{p-1}(X/k,G)$ be the category of nilpotent flat $G$-bundles on $X$ of exponent $\leq p-1$. Furthermore, if $G={\rm GL}_{n}(k)$, we use the notation ${\rm MIC}_{p-1}(X/k)$ for the category of nilpotent flat $G$-bundles of exponent $\leq p-1$.
Remark 2.6.
The exponent of nilpotency for $G$-Higgs bundles and flat $G$-bundles considered in this paper is only defined for the integer $p$. Now we consider a special case that the exponent is equal to zero. In this case, a nilpotent Higgs field is trivial. Thus, a nilpotent Higgs bundle of exponent $\leq 0$ is exactly a $G$-bundle, and then the category ${\rm HIG}_{0}(X/k,G)$ of nilpotent $G$-Higgs bundles of exponent $\leq 0$ is exactly the category of $G$-bundles. Similarly, a nilpotent flat $G$-bundle of exponent $\leq 0$ implies that it has trivial $p$-curvature. Thus, the category ${\rm MIC}_{0}(X/k,G)$ of nilpotent flat $G$-bundles of exponent $\leq 1$ is the category of flat $G$-bundles with trivial $p$-curvatures.
2.3. Stability Conditions
In this subsection, we suppose that $X$ is a smooth projective variety and we fix an ample line bundle $H$ on $X$. Then the degree of a vector bundle $E$ on $X$ is given with respect to $H$ and denote it by $\deg_{H}E$. If there is no ambiguity, we use the notation $\deg E$ and omit the subscript $H$.
G-bundles
Ramanathan studied stability conditions of principal bundles on curves in characteristic zero [Ram75, Ram96a, Ram96b]. We generalize his definition to smooth projective varieties in characteristic $p$, and we call it the $R$-stability condition. Let $P$ be a parabolic subgroup of $G$ together with a reduction of structure group $\sigma:X\rightarrow E/P$. Then, we get a $P$-bundle $E_{\sigma}$ in the following way
[…]
Given a character $\chi:P\rightarrow\mathbb{G}_{m}$, we obtain a line bundle $\chi_{*}E_{\sigma}$ on $X$ and denote it by $L(\sigma,\chi)$.
Definition 2.7 ($R$-stability condition).
A $G$-bundle $E$ is $R$-semistable (resp. $R$-stable), if
  • for any proper parabolic subgroup $P\subseteq G$;
  • for any antidominant character $\chi:P\rightarrow\mathbb{G}_{m}$ acting trivially on the center (we refer the reader to [Ram75] for the definition of anti-dominant characters);
  • for any reduction of structure group $\sigma:X\rightarrow E/P$,
we have
[…]
G-Higgs bundles
Now we consider the $R$-stability condition of $G$-Higgs bundles and we refer the reader to [KSZ24] for the case of curves. Let $\sigma:X\rightarrow E/P$ be a reduction of structure group. We obtain a $P$-bundle $E_{\sigma}$ together with a map $E_{\sigma}\rightarrow E$. Therefore, we obtain a natural map
$E_{\sigma}(\mathfrak{p})\rightarrow E({\mathfrak{g}}).$
We say that $\sigma$ is $\theta$-compatible if there is a lifting $\theta_{\sigma}:X\rightarrow E_{\sigma}(\mathfrak{p})\otimes\Omega_{X}$ such that the diagram commutes
[…]
Remark 2.8.
Recall that a Higgs bundle $(E,\theta)$ is semistable (resp. stable), if for all $\theta$-invariant subbundles $F$, we have
$\frac{\deg F}{{\rm rk}F}\leq\frac{\deg E}{{\rm rk}E}\quad(\text{resp. }<).$. < ) .
Note that we take “$\theta$-invariant subbundles” to define the stability condition of a Higgs bundle. In the case of $G$-Higgs bundles, we choose $\theta$-compatible reductions for the same reason (see [KSZ24, §4]).
Definition 2.9.
A $G$-Higgs bundle $(E,\theta)$ is $R$-semistable (resp. $R$-stable), if
  • for any proper parabolic subgroup $P\subseteq G$;
  • for any antidominant character $\chi:P\rightarrow\mathbb{G}_{m}$ acting trivially on the center;
  • for any $\theta$-compatible reduction of structure group $\sigma:X\rightarrow E/P$,
we have
[…]
Denote by ${\rm HIG}^{s(s)}_{p-1}(X/k,G)$ the category of $R$-(semi)stable nilpotent $G$-Higgs bundles of exponent $\leq p-1$, which is a full subcategory of ${\rm HIG}_{p-1}(X/k,G)$.
Flat G-bundles
Let $(V,\nabla)$ be a flat $G$-bundle. Let $\varsigma$ be a reduction of structure group, and then we have a map $V_{\varsigma}\rightarrow V$ (also $\mathcal{O}_{V}\rightarrow\mathcal{O}_{V_{\varsigma}}$). We say that $\varsigma$ is $\nabla$-compatible if the connection $\nabla$ induces a unique connection $\nabla_{\varsigma}:\mathcal{O}_{V_{\varsigma}}\rightarrow\mathcal{O}_{V_{\varsigma}}\otimes\Omega_{X}$ such that the diagram commutes
[…]
When $V$ is a trivial $G$-bundle (or consider $V$ locally), $\nabla$ can be regarded as a section of $V(\mathfrak{g})\otimes\Omega_{X}$ (Example 2.3). In this case, the $\nabla$-compatibility can be defined in the same way as that of $G$-Higgs bundles. More precisely, a $\nabla$-compatible reduction $\varsigma:X\rightarrow V/P$ is equivalent to the condition that there exists a connection $\nabla_{\varsigma}$ of $V_{\varsigma}$ such that the diagram commutes
[…]
Definition 2.10.
A flat $G$-bundle $(V,\nabla)$ is $R$-semistable (resp. $R$-stable), if
  • for any proper parabolic subgroup $P\subseteq G$;
  • for any antidominant character $\chi:P\rightarrow\mathbb{G}_{m}$ acting trivially on the center;
  • for any $\nabla$-compatible reduction of structure group $\varsigma:X\rightarrow V/P$,
we have
[…]
Let ${\rm MIC}^{(s)s}_{p-1}(X/k,G)$ be the category of $R$-(semi)stable nilpotent flat $G$-bundles of exponent $\leq p-1$.
3. Nilpotent $G$-Higgs Bundles and Nilpotent Flat $G$-Bundles
The goal of this section is proving the equivalence of categories under the $W_{2}(k)$-liftable condition:
Theorem 3.1.
Suppose that $X$ is $W_{2}(k)$-liftable. The categories ${\rm HIG}_{p-1}(X^{\prime}/k,G)$ and ${\rm MIC}_{p-1}(X/k,G)$ are equivalent.
To prove the theorem, we construct inverse Cartier functor in §3.1
$C^{-1}_{\rm exp}:{\rm HIG}_{p-1}(X^{\prime}/k,G)\rightarrow{\rm MIC}_{p-1}(X/k,G),$
and Cartier functor in §3.2
$C_{\rm exp}:{\rm MIC}_{p-1}(X/k,G)\rightarrow{\rm HIG}_{p-1}(X^{\prime}/k,G)$
via an approach called exponential twisting introduced in [LSZ15]. These two functors are inverse to each other, and then induces the equivalence of categories. Furthermore, the equivalence holds under stability conditions. In other words, categories ${\rm HIG}^{(s)s}_{p-1}(X^{\prime}/k,G)$ and ${\rm MIC}^{(s)s}_{p-1}(X/k,G)$ are equivalent (Theorem 3.6). The key point to generalize Lan–Sheng–Zuo’s approach to principal bundles is to give an appropriate definition of the exponentiation in the case of principal bundles, which has been done by a series work (see [Ser94, Sei00, McN02, Sob15] for instance) and reviewed in §2. Before we move to the Cartier and inverse Cartier transform, we first review two classical results: classical Cartier descent and Deligne-Illusie’s lemma, which will be used frequently in this section.
Classical Cartier descent
The classical Cartier descent is given as follows:
Theorem 3.2 (Theorem 5.1 in [Kat70]).
There is an equivalence of categories between ${\rm HIG}_{0}(X^{\prime}/k)$ and ${\rm MIC}_{0}(X/k)$.
Given $(E,0)\in{\rm HIG}_{0}(X^{\prime}/k)$, Katz constructed a canonical connection $\nabla_{can}$ on $F^{*}_{X/k}E$, which induces a functor ${\rm HIG}_{0}(X^{\prime}/k)\rightarrow{\rm MIC}_{0}(X/k)$, and prove that the categories are equivalent. Chen–Zhu extended this result to principal bundles:
Corollary 3.3 (Example A.5 in [CZ15]).
Let $\mathcal{G}^{\prime}$ be a smooth affine group scheme over $X^{\prime}$ and define $\mathcal{G}:=F^{*}_{1}\mathcal{G}^{\prime}$. Then we have an equivalence of categories between the category of $\mathcal{G}^{\prime}$-torsors on $X^{\prime}$ and the category of $\mathcal{G}$-local systems with trivial $p$-curvature. As a special case, when $\mathcal{G}^{\prime}=X^{\prime}\times G$, then the categories ${\rm HIG}_{0}(X^{\prime}/k,G)$ and ${\rm MIC}_{0}(X/k,G)$ are equivalent.
A Lemma by Deligne-Illusie
We review a lemma by Deligne–Illusie and we refer the reader to [DI87] and [LSZ15] for more details. Let $W_{2}(k)$ be the ring of Witt vectors of length two. We fix a $W_{2}(k)$-lifting $X_{2}$ of $X$, and choose an open affine cover $\{U_{2,i}\}_{i\in I}$ of $X_{2}$, where $I$ is the index set. Let $F_{2,i}:U_{2,i}\rightarrow U^{\prime}_{2,i}$ be a Frobenius lifting, which ${\rm mod}\,p$ is the relative Frobenius morphism $F_{1,i}:U_{i}\rightarrow U^{\prime}_{i}$, and we use the notation $F_{1}:=F_{1,i}$ for convenience. We have a natural morphism
$\zeta_{i}:=\frac{F^{*}_{2,i}}{[p]}:F^{*}_{1}\Omega_{U^{\prime}_{i}}\rightarrow\Omega_{U_{i}}.$
Given $i,j,k\in I$, define $U_{ij}:=U_{i}\cap U_{j}$ and $U_{ijk}:=U_{i}\cap U_{j}\cap U_{k}$ to be the intersections.
Lemma 3.4 (§2 in [DI87] or Lemma 2.1 in [LSZ15]).
There exist homomorphisms $h_{ij}:F_{1}^{*}\Omega_{U^{\prime}_{ij}}\rightarrow\mathcal{O}_{U_{ij}}$ such that
  1. (1)
    $\zeta_{i}-\zeta_{j}=dh_{ij}$ over $U_{ij}$,
  2. (2)
    $h_{ij}+h_{jk}=h_{ik}$ over $U_{ijk}$.
3.1. Inverse Cartier Transform
In this section, we will show that given a nilpotent $G$-Higgs bundle $(E,\theta)$, it can be associated with a nilpotent flat $G$-bundle. Furthermore, this construction induces a functor
$C^{-1}_{\rm exp}:{\rm HIG}_{p-1}(X^{\prime}/k,G)\rightarrow{\rm MIC}_{p-1}(X/k,G),$
We fix some notations first. Let $\mathcal{U}=\{U_{i}\}_{i\in I}$ be a covering of open affine subsets of $X$ given in Lemma 3.4. Given a $G$-Higgs bundle $(E,\theta)$ on $X^{\prime}$, denote by $E_{i}:=E|_{U_{i}}$ the restriction of $E$ to the open set $U_{i}$. We can define the above notations on the intersections $U_{ij}$ similarly. For convenience, we use the same notation $F_{1}$ for relative Frobenius maps of $X$, $U_{i}$ and etc. for simplicity.
Let $(E,\theta)$ be a $G$-Higgs bundle in ${\rm HIG}_{p-1}(X^{\prime}/k,G)$. The element $h_{ij}(F_{1}^{*}\theta)$ is a section of $(F_{1}^{*}E_{ij})(\mathfrak{g})$. Since $\theta$ is nilpotent of exponent $\leq p-1$, the element
$G_{ij}:=\overline{\exp}(h_{ij}(F_{1}^{*}\theta))$
is well-defined by Theorem 2.5 and induces an automorphism of $F^{*}_{1}E_{ij}$. Since $\theta$ is integrable, we have
$\overline{\exp}(h_{ij}(F^{*}_{1}\theta))\cdot\overline{\exp}(h_{jk}(F^{*}_{1}\theta))=\overline{\exp}((h_{ij}+h_{jk})(F^{*}_{1}\theta)).$
By Lemma 3.4, the data $\{G_{ij}\}_{i,j\in I}$ satisfies the cocycle condition
(3.1) $G_{ij}\cdot G_{jk}=G_{ik}.$
By gluing $\{V_{i}:=F^{*}_{1}E_{i}\}_{i\in I}$ via $G_{ij}$, we obtain a $G$-bundle $V$ on $X$.
Next, we construct a connection on $V$. Denote by $\nabla_{can,i}$ the integrable connection on $F^{*}_{1}E_{i}$ given by Corollary 3.3, and we define a new connection
(3.2) $\nabla_{i}=\nabla_{can,i}+\zeta_{i}(F^{*}_{1}\theta)$
on $F^{*}_{1}E_{i}$. We have to show that the local data $\{\nabla_{i}\}_{i\in I}$ can be glued together via $\{G_{ij}\}_{i,j\in I}$, i.e.
(3.3) $G_{ij}\circ\nabla_{j}=\nabla_{i},$
where the action is the gauge action. It is equivalent to show
$dG_{ij}\cdot G_{ij}^{-1}+{\rm Ad}(G_{ij})(\zeta_{j}(F^{*}_{1}\theta))=\zeta_{i}(F^{*}_{1}\theta).$
The first term $G_{ij}^{-1}\cdot dG_{ij}$ is $dh_{ij}(F_{1}^{*}\theta)$. For the second term ${\rm Ad}(G_{ij})(\zeta_{j}(F^{*}_{1}\theta))$, we have
${\rm Ad}(G_{ij})(\zeta_{j}(F^{*}_{1}\theta))={\rm Ad}(\overline{\exp}(h_{ij}(F_{1}^{*}\theta)))(\zeta_{j}(F^{*}_{1}\theta))=\zeta_{j}(F^{*}_{1}\theta)$
by the second property of the exponential map (Definition 2.4). Summing up, we have
$dG_{ij}\cdot G_{ij}^{-1}+{\rm Ad}(G_{ij})(\zeta_{j}(F^{*}_{1}\theta))=dh_{ij}(F_{1}^{*}\theta)+\zeta_{j}(F^{*}_{1}\theta).$
Applying Lemma 3.4 again, we get the desired equality and obtain a $G$-connection $\nabla$ on $V$.
Finally, we will prove that the $G$-connection $\nabla$ is integrable and nilpotent of exponent $\leq p-1$. Since either integrability and nilpotency is a local property, we only have to check them on local charts $U_{i}$. Then the proof of the integrability is exactly the same as [LSZ15, §2.2 Step 3]. For the nilpotency, the nilpotency of the $p$-curvature of $\nabla_{i}$ is equivalent to that of $F^{*}_{1}\theta_{i}$, and thus, $\nabla_{i}$ is also nilpotent of exponent $\leq p-1$.
In conclusion, we obtain a functor
$C^{-1}_{\rm exp}:{\rm HIG}_{p-1}(X^{\prime}/k,G)\rightarrow{\rm MIC}_{p-1}(X/k,G).$
3.2. Cartier Transform
We consider the other direction: given a flat $G$-bundle $(V,\nabla)\in{\rm MIC}_{p-1}(X/k,G)$, we want to construct a $G$-Higgs bundle $(E,\theta)\in{\rm HIG}_{p-1}(X^{\prime}/k,G)$, which will induce a functor
$C_{\rm exp}:{\rm MIC}_{p-1}(X/k,G)\rightarrow{\rm HIG}_{p-1}(X^{\prime}/k,G).$
Given a pair $(V,\nabla)\in{\rm MIC}_{p-1}(X/k,G)$, denote by $\psi$ the $p$-curvature of $\nabla$. Note that $\psi$ is nilpotent of exponent $\leq p$ by definition. We use the same notation for coverings as in §3.1, and let $V_{i}:=V|_{U_{i}}$, $\psi_{i}:=\psi|_{U_{i}}$. Then, we define a new connection $\nabla^{\prime}_{i}$ on $V_{i}$ as
$\nabla^{\prime}_{i}:=\nabla_{i}+\zeta_{i}(\psi).$
Following the same argument as [LSZ15, §2.3, Step 3], the $p$-curvature of $\nabla^{\prime}_{i}$ is trivial. Moreover, since the $p$-curvature $\psi_{i}$ is horizontal with respect to $\nabla_{i}$, it is also horizontal with respect to $\nabla^{\prime}_{i}$. Now we want to glue $\{(V_{i},\nabla^{\prime}_{i},\psi_{i})\}_{i\in I}$ to obtain a new flat $G$-bundle $(V^{\prime},\nabla^{\prime})$, of which the $p$-curvature is trivial, together with a $F$-Higgs field $\psi$ horizontal with respect to $\nabla^{\prime}$. We define
$J_{ij}:=\overline{\exp}(h_{ij}(\psi)),$
which is an automorphism of $V_{ij}$. With the same argument as formula (3.1), we have
$J_{ij}\cdot J_{jk}=J_{ik},\quad J_{ij}\circ\psi_{j}=\psi_{i},$
where $J_{ij}\circ\psi_{j}$ is the adjoint action, and then we obtain a new $G$-bundle $V^{\prime}$ and a new section $\psi^{\prime}$ of $V^{\prime}(\mathfrak{g})\otimes F^{*}_{1}\Omega_{X^{\prime}}$, which is regarded as a $F$-Higgs field. A similar calculation of formula (3.3) gives
$\quad J_{ij}\circ\nabla^{\prime}_{j}=\nabla^{\prime}_{i},$
and we get an integrable $G$-connection $\nabla^{\prime}$ on $V^{\prime}$. In conclusion, we get a triple $(V^{\prime},\nabla^{\prime},\psi^{\prime})$, where $(V^{\prime},\nabla^{\prime})$ is a flat $G$-bundle with trivial $p$-curvature and $\psi^{\prime}$ is a $F$-Higgs field on $V^{\prime}$. Since $\psi^{\prime}$ is horizontal with respect to $\nabla^{\prime}$, of which the $p$-curvature is zero, $(V^{\prime},\nabla^{\prime},\psi^{\prime})$ descends to a $G$-Higgs bundle $(E,\theta)$ by Corollary 3.3. Therefore, we obtain a functor
$C_{\rm exp}:{\rm MIC}_{p-1}(X/k,G)\rightarrow{\rm HIG}_{p-1}(X^{\prime}/k,G).$
as expected.
3.3. Stability Condition
Suppose that $X$ is smooth and projective. We fix an ample line bundle $H$ on $X$ and define $H^{\prime}:=\pi_{X/k}^{*}H$, which is also an ample line bundle on $X^{\prime}$. Furthermore, $H^{p}\cong F_{X/k}^{*}H^{\prime}=F_{X}^{*}H$. Note that the stability condition of Higgs bundles on $X^{\prime}$ is given by the ample line bundle $H^{\prime}$ while the stability condition of flat bundles on $X$ is given by $H$. In this subsection, we improve the equivalence given in Corollary 3.3 by adding the information of stability conditions:
Proposition 3.5.
There is an equivalence of categories between ${\rm HIG}^{(s)s}_{0}(X^{\prime}/k,G)$ and ${\rm MIC}^{(s)s}_{0}(X/k,G)$.
Proof.
Let $E$ be a $G$-bundle on $X^{\prime}$, and denote by $(V,\nabla_{can})$ the corresponding flat $G$-bundle on $X$ with trivial $p$-curvature by Corollay 3.3, where $V=F^{*}_{1}E$. Let $P$ be a parabolic subgroup of $G$. To prove the equivalence of stability conditions, it is enough to prove the following two statements:
  1. (1)
    there is a one-to-one correspondence between reductions of structure group $\sigma:X^{\prime}\rightarrow E/P$ and $\nabla_{can}$-compatible reductions of structure group $\varsigma:X\rightarrow V/P$;
  2. (2)
    we have
    $p\deg_{H}L(\varsigma,\chi)=\deg_{H^{\prime}}L(\sigma,\chi),$
    where reductions $\sigma$ and $\varsigma$ are given by the first statement and $\chi:P\rightarrow\mathbb{G}_{m}$ is a character.
For the first statement, given a reduction of structure group $\sigma:X^{\prime}\rightarrow E/P$, we obtain a reduction $\varsigma:X\rightarrow V/P$ by pullback under the relative Frobenius morphism $F_{1}$. In other words, the morphism $E_{\sigma}\rightarrow E$ induces a morphism $(V_{\varsigma},\nabla_{can,\varsigma})\rightarrow(V,\nabla_{can})$ of flat $G$-bundles by the classical Cartier descent (Theorem 3.2), where $V_{\varsigma}=F^{*}_{1}E_{\sigma}$ and $\nabla_{can,\varsigma}$ is the canonical connection on $V_{\varsigma}$ given by Cartier descent. Therefore, the diagram
${\mathcal{O}_{V}}$${\mathcal{O}_{V}\otimes\Omega_{X}}$${\mathcal{O}_{V_{\varsigma}}}$${\mathcal{O}_{V_{\varsigma}}\otimes\Omega_{X}}$$\scriptstyle{\nabla_{can}}$$\scriptstyle{\nabla_{can,\varsigma}}$
commutes. This shows that $\varsigma$ is $\nabla_{can}$-compatible. By the classical Cartier descent, it is easy to see that this is a one-to-one correspondence. The second statement is clear because $H^{p}\cong F^{*}_{1}H^{\prime}$ and $V_{\varsigma}=F^{*}_{1}E_{\sigma}$. ∎
Now we consider the case of nilpotency of exponent $\leq p-1$.
Theorem 3.6.
The categories ${\rm HIG}^{(s)s}_{p-1}(X^{\prime}/k,G)$ and ${\rm MIC}^{(s)s}_{p-1}(X/k,G)$ are equivalent.
Proof.
Let $(E,\theta)\in{\rm HIG}_{p-1}(X^{\prime}/k,G)$ be a $G$-Higgs bundle, and denote by $(V,\nabla)\in{\rm MIC}_{p-1}(K/k,G)$ the corresponding flat $G$-bundle under the equivalence given by Theorem 3.1. To prove the theorem, it is enough to show that $\theta$-compatible reductions $\sigma$ of $E$ are in one-to-one correspondence with $\nabla$-compatible reductions $\varsigma$ of $V$.
Let $\sigma:X^{\prime}\rightarrow E/P$ be a $\theta$-compatible reduction. Denote by $\theta_{\sigma}:X^{\prime}\rightarrow E_{\sigma}(\mathfrak{p})\otimes\Omega_{X^{\prime}}$ the lifting of $\theta$ via $\sigma$. We work on local charts $U_{i}$ first and construct a $\nabla_{i}$-compatible reduction $\varsigma_{i}$, where $\nabla_{i}:=\nabla|_{U_{i}}$, and then we will show that the local data $\varsigma_{i}$ can be glued together. We use the same notation as we did in §3.1 and §3.2. Denote by $\sigma_{i}=\sigma|_{U_{i}}$ the restriction to $U_{i}$. In §3.1, we construct $(V_{i},\nabla_{i})$ on $U_{i}$, where $V_{i}=F^{*}_{1}E_{i}$ and $\nabla_{i}=\nabla_{can}+\zeta_{i}(F^{*}_{1}\theta)$. Let $\varsigma_{i}:U_{i}\rightarrow V_{i}/P$ be the corresponding reduction of $\sigma_{i}$ given in Proposition 3.5, and then $\varsigma_{i}$ is compatible with $\nabla_{can}$. Since $\sigma_{i}$ is compatible with $\theta_{i}$, the reduction $\varsigma_{i}$ is compatible with $F^{*}_{1}\theta_{i}$ and thus compatible with $\zeta_{i}(F^{*}_{1}\theta)$. In conclusion, the reduction $\varsigma_{i}$ is compatible with $\nabla_{i}$, and we denote by $\nabla_{\varsigma,i}$ the induced $P$-connection on $V_{\varsigma,i}$. Then, we have the following diagrams on $U_{i}$.
[…]
Moreover, we obtain a flat $P$-bundle $(V_{\varsigma},\nabla_{\varsigma})$ by gluing local data $(V_{\varsigma,i},\nabla_{\varsigma,i})$ via
$G_{\sigma,ij}:=\overline{\exp}(h_{ij}(F^{*}_{1}(\theta_{\sigma})))$
with the same argument as in §3.1.
Now we will show that $\varsigma_{i}$ can be glued together via $G_{ij}$, and it is equivalent to prove that the morphisms $\{V_{\varsigma,i}\rightarrow V_{i}\}_{i\in I}$ can be glued together. For convenience, we use the same notation $\varsigma_{i}$ for the induced morphism $V_{\varsigma,i}\rightarrow V_{i}$, and we have to show that the following diagram commutes.
[…]
By the compatibility of $\nabla_{i}$ and $\nabla_{j}$, we have
$\nabla_{j}=\varsigma_{j}\circ\nabla_{\varsigma,j},\quad\nabla_{i}=\varsigma_{i}\circ\nabla_{\varsigma,i}.$
Since $\{\nabla_{\varsigma,i},\,i\in I\}$ and $\{\nabla_{i},\,i\in I\}$ can be glued together via $G_{\sigma,ij}$ and $G_{ij}$ respectively, we have
$G_{\sigma,ij}\circ\nabla_{\varsigma,j}=\nabla_{\varsigma,i},\quad G_{ij}\circ\nabla_{j}=\nabla_{i}.$
Thus, we obtain the desired property
$G_{ij}\circ\varsigma_{j}=\varsigma_{i}\circ G_{\sigma,ij}$
by the following diagram.
[…]
In conclusion, we obtain a $\nabla$-compatible reduction $\varsigma:X\rightarrow V/P$.
The other direction is proved in a similar way. Let $\varsigma:X\rightarrow V/P$ be a $\nabla$-compatible reduction. Clearly, $\varsigma$ is also compatible with the $p$-curvature $\psi$. Denote by $\nabla_{\varsigma}$ (resp. $\psi_{\varsigma}$) the induced connection ($p$-curvature) by $\nabla$. Recall that we defined a triple $(V_{i},\nabla^{\prime}_{i},\psi_{i})$ on each $U_{i}$, where
$V_{i}=V|_{U_{i}},\quad\nabla^{\prime}_{i}=\nabla_{i}+\zeta_{i}(\psi),\quad\psi_{i}=\psi|_{U_{i}}.$
Since $\varsigma$ is compatible with $\nabla$ and $\psi$, the restriction $\varsigma_{i}:=\varsigma|_{U_{i}}$ is also compatible with $\nabla_{i}$ and $\psi_{i}$ (thus compatible with $\nabla^{\prime}_{i}$). Let $V_{\varsigma,i}$ be the restriction of the $P$-bundle $V_{\varsigma,i}$ to $U_{i}$, and denote by $\nabla_{\varsigma,i}$ the induced connection on $V_{\varsigma,i}$. Abusing the notation, we use the same notation $\varsigma_{i}:V_{\varsigma,i}\rightarrow V_{i}$ for the induced morphism. Locally, we have the following diagram
[…]
where
$J_{\varsigma,ij}:=\overline{\exp}(h_{ij}(\psi_{\varsigma})),$
such that
  • $\psi$-compatibility:
    $\psi_{j}=\varsigma_{j}\circ\psi_{\varsigma,j},\quad\psi_{i}=\varsigma_{i}\circ\psi_{\varsigma,i}.$
  • gluing property:
    $J_{\varsigma,ij}\circ\psi_{\varsigma,j}=\psi_{\varsigma,i},\quad J_{ij}\circ\psi_{j}=\psi_{i}.$
Therefore,
$J_{ij}\circ\varsigma_{j}=\varsigma_{i}\circ J_{\varsigma,ij},$
and then the local data $\{\varsigma_{i}\}_{i\in I}$ can be glued via $J_{ij}$, and denote it by $\varsigma^{\prime}$. Clearly, the reduction $\varsigma^{\prime}$ is compatible with both $\nabla^{\prime}$ and $\psi^{\prime}$. By Proposition 3.5, $(V^{\prime},\nabla^{\prime})$ descent to a $G$-bundle $E$ on $X^{\prime}$. Thus, $\varsigma^{\prime}$ descents to a reduction $\sigma:X^{\prime}\rightarrow E/P$. Since $\varsigma^{\prime}$ is $\psi^{\prime}$-compatible, the reduction $\sigma$ is $\theta$-compatible. This finishes the proof. ∎
4. Logarithmic G-Higgs bundles and Flat G-bundles on Root Stacks
In this section, we are mostly interested in a special classes of smooth Deligne–Mumford stacks $\mathscr{X}$ considered and discussed in [MO05, Cad07, Sim11], which are called root stacks in this paper (Theorem 4.3). We define logarithmic nilpotent $G$-Higgs bundles and logarithmic nilpotent flat $G$-bundles on $\mathscr{X}$ of exponent $\leq p-1$ and prove an equivalence of categories between ${\rm HIG}_{p-1}(\mathscr{X}^{\prime}_{\rm log}/k,G)$ and ${\rm MIC}_{p-1}(\mathscr{X}_{\rm log}/k,G)$ under a $W_{2}$-lifting condition (Theorem 4.12). Then, we follow Borne and Nironi’s approach to give the stability condition [Bor07, Nir09], and prove the equivalence of categories under stability conditions (Theorem 4.16).
4.1. G-Higgs Bundles and Flat G-bundles on DM stacks
Recall that a Deligne–Mumford stack over $k$ is an algebraic stack such that there exists a surjective étale morphism $Y\rightarrow\mathscr{X}$, where $Y$ is a scheme. It is well-known that $\mathscr{X}\cong[Y/R]$, where $R=Y\times_{\mathscr{X}}Y$ is an algebraic space and we have two étale morphisms
$p_{i}:R=Y\times_{\mathscr{X}}Y\rightarrow Y$
for $i=1,2$. For convenience, we suppose that $R$ is a scheme. In fact, we can apply all arguments in this section to algebraic spaces first, and then give the proof for Deligne–Mumford stacks. Let $F_{\mathscr{X}/k}:\mathscr{X}\rightarrow\mathscr{X}^{\prime}$ be the relative Frobenius morphism, for which we still use the notation $F_{1}:=F_{\mathscr{X}/k}$ for simplicity, and we suppose that $\mathscr{X}$ is $W_{2}(k)$-liftable. Here is an equivalent description under the isomorphism $\mathscr{X}\cong[Y/R]$. The relative Frobenius morphism $F_{\mathscr{X}/k}$ is induced from those on $Y$ and $R$, and $\mathscr{X}$ is $W_{2}(k)$-liftable is equivalent to say that $Y$ and $R$ are $W_{2}(k)$-liftable and the following diagrams commutes
[…]
The stack $\mathscr{X}$ is called tame if the induced functor
$\pi_{*}:{\rm QCoh}(\mathscr{X})\rightarrow{\rm QCoh}(X)$
is exact, where ${\rm QCoh}$ is the category of quasi-coherent sheaves. We want to remind the reader that the functor $\pi_{*}$ is always exact in the case of characteristic zero. In positive characteristic, there are several equivalent conditions of tameness [AOV08, Theorem 3.2].
Suppose that $\mathscr{X}$ is tame, smooth and $W_{2}(k)$-liftable with coarse moduli space $\pi:\mathscr{X}\rightarrow X$, where $X$ is a smooth variety over $k$. Denote by $\Omega_{\mathscr{X}/k}$ the relative cotangent sheaf, and if there is no confusion, we use the notation $\Omega_{\mathscr{X}}$ for simplicity. Suppose that $\mathscr{X}\cong[Y/R]$, and then a $G$-bundle $\mathscr{E}$ on $\mathscr{X}$ can be regarded as a pair $(E,\tau)$, where $E$ is a $G$-bundle on $Y$ and $\tau:p_{1}^{*}E\xrightarrow{\cong}p_{2}^{*}E$ is an isomorphism. In the same way, a $G$-Higgs bundle $(\mathscr{E},\theta)$ on $\mathscr{X}$ is regarded as a triple $(E,\theta,\tau)$ on $Y$, where $(E,\theta)$ is a Higgs bundle on $Y$ and $\tau:p_{1}^{*}(E,\theta)\rightarrow p_{2}^{*}(E,\theta)$ is an isomorphism, and a similar description can be given for flat $G$-bundles $(\mathscr{V},\nabla)$ on $\mathscr{X}$. Therefore, the nilpotency for $G$-Higgs bundles and flat $G$-bundles on $\mathscr{X}$ is defined as we did on varieties in §2.2. Define the follow categories:
  • ${\rm HIG}_{0}(\mathscr{X}^{\prime}/k,G)$ is the category of $G$-bundles on $\mathscr{X}^{\prime}$,
  • ${\rm MIC}_{0}(\mathscr{X}/k,G)$ is the category of flat $G$-bundles on $\mathscr{X}$ with trivial $p$-curvature,
  • ${\rm HIG}_{p-1}(\mathscr{X}^{\prime}/k,G)$ is the category of nilpotent $G$-Higgs bundles on $\mathscr{X}^{\prime}$ of exponent $\leq p-1$,
  • ${\rm MIC}_{p-1}(\mathscr{X}/k,G)$ is the category of nilpotent flat $G$-bundles on $\mathscr{X}$ of exponent $\leq p-1$.
Moreover, in the rest of the paper, we only give the definition of categories of objects with nilpotent exponent $\leq p-1$ and omit the deifnition for objects with nilpotent exponent $\leq 0$ for simplicity.
Proposition 4.1.
We have an equivalence of categories between ${\rm HIG}_{0}(\mathscr{X}^{\prime}/k,G)$ and ${\rm MIC}_{0}(\mathscr{X}/k,G)$.
Proof.
With the same setup as above, we prove the equivalence via the isomorphism $\mathscr{X}\cong[Y/R]$. We have the following commutative diagram,
[…]
where $F_{1}$ is the relative Frobenius morphism. Given a $G$-bundle $\mathscr{E}$ on $\mathscr{X}^{\prime}$, it can be regarded as a pair $(E,\tau)$, where $E$ is a $G$-bundle on $Y^{\prime}$ and $\tau:p_{1}^{*}E\cong p_{2}^{*}E$ is an isomorphism of $G$-bundles on $R^{\prime}$. Since $F_{1}\circ p_{i}=p_{i}\circ F_{1}$, we have $F_{1}^{*}p_{i}^{*}E\cong p_{i}^{*}F_{1}^{*}E$. By Corollary 3.3, the $G$-bundle $E$ corresponds to a flat $G$-bundle $(V,\nabla_{can})$ on $Y$, where $V=F^{*}_{1}E$. On the other hand, the $G$-bundle $p_{i}^{*}E$ on $R^{\prime}$ corresponds to a flat $G$-bundle on $R$, which is isomorphic to $p_{i}^{*}(V,\nabla_{can})$ by the isomorphism $F_{1}^{*}p_{i}^{*}E\cong p_{i}^{*}F_{1}^{*}E$ and uniqueness of the canonical connection. Thus, we obtain an isomorphism $p_{1}^{*}(V,\nabla_{can})\cong p_{2}^{*}(V,\nabla_{can})$ induced from $\tau:p_{1}^{*}E\rightarrow p_{2}^{*}E$, and denote it by $F_{1}^{*}\tau$. The data $((V,\nabla_{can}),F^{*}\tau)$ gives a flat $G$-bundle $(\mathscr{V},\nabla_{can})$ with trivial $p$-curvature on $\mathscr{X}$. The other direction can be proved similarly. ∎
Theorem 4.2.
The categories ${\rm HIG}_{p-1}(\mathscr{X}^{\prime}/k,G)$ and ${\rm MIC}_{p-1}(\mathscr{X}/k,G)$ are equivalent.
Proof.
The proof is similar to that of Proposition 4.1, and we still work on $\mathscr{X}\cong[Y/R]$. By Theorem 3.1, we have the following equivalences of categories:
  • ${\rm HIG}_{p-1}(Y^{\prime}/k,G)$ and ${\rm MIC}_{p-1}(Y/k,G)$;
  • ${\rm HIG}_{p-1}(R^{\prime}/k,G)$ and ${\rm MIC}_{p-1}(R/k,G)$.
Given a $G$-Higgs bundle $(\mathscr{E},\theta)$ on $\mathscr{X}^{\prime}$, we regard it as a $G$-Higgs bundle $(E,\theta)$ on $Y^{\prime}$ together with an isomorphism $\tau$. The first equivalence of categories gives a flat $G$-bundle $(V,\nabla)\in{\rm MIC}_{p-1}(Y/k,G)$, and then we obtain flat $G$-bundles $p_{i}^{*}(V,\nabla)\in{\rm MIC}_{p}(R/k,G)$ for $i=1,2$. Via étale morphisms $p_{i}$ for $i=1,2$, we obtain $G$-Higgs bundles on $R^{\prime}$, which are isomorphic to $p_{i}^{*}(E,\theta)$ by the commutative diagram in Proposition 4.1. Therefore, the isomorphism $\tau$ induces an isomorphism $p_{1}^{*}(V,\nabla)\cong p_{2}^{*}(V,\nabla)$, and denote it by $F^{*}\tau$. Thus, the data $((V,\nabla),F^{*}\tau)$ gives a flat $G$-bundle in ${\rm MIC}_{p-1}(\mathscr{X}/k,G)$. The proof of the other direction is given in the same way. ∎
4.2. G-bundles on Root Stacks
We consider a special class of tame smooth Deligne–Mumford stacks given in [MO05, Cad07], which are called Cadman–Vistoli root stacks in [Sim11]. We are mostly interested in these root stacks in the rest of this paper. Let $X$ be a smooth variety with a reduced normal crossing divisor $D=\sum\limits_{i=1}^{s}D_{i}$. We equip each irreducible component $D_{i}$ with a positive integer $d_{i}$ coprime to $p$. Denote by $\boldsymbol{d}=\{d_{i},1\leq i\leq s\}$ the collection of these integers.
Theorem 4.3 (Theorem 4.1 in [MO05]).
Given the data $(X,D,\boldsymbol{d})$ above, there exists a tame smooth Deligne–Mumford stack $\pi:\mathscr{X}\rightarrow X$ together with a reduced normal crossing divisor $\widetilde{D}=\sum\limits_{i=1}^{s}\widetilde{D}_{i}\subseteq\mathscr{X}$ such that
  • the coarse moduli space of $\mathscr{X}$ is $X$,
  • the map $\pi$ is finite flat and induces an isomorphism over $X\backslash D$,
  • the pullback $\pi^{*}\mathcal{O}_{X}(-D_{i})$ is equal to $\mathcal{O}_{\mathscr{X}}(-d_{i}\widetilde{D_{i}})$.
We use the notation $\mathscr{X}_{(X,D,\boldsymbol{d})}$ for this stack and call it a root stack in this paper.
Example 4.4.
The example has been discussed in [MO05, Lemma 4.3], and we include it here because it is very helpful to understand the local structure around a point $\widetilde{x}\in\widetilde{D}\subseteq\mathscr{X}$. Let $X={\rm Spec}\,k[z_{1},\dots,z_{d}]$ and $D_{i}=Z(z_{i})$ for $1\leq i\leq s$, and we suppose that $s\leq d$. Given a collection of positive integers $\boldsymbol{d}=\{d_{i},1\leq i\leq s\}$, we define a natural morphism
$k[z_{1},\dots,z_{d}]\longrightarrow k[w_{1},\dots,w_{d}]$
such that
$z_{i}\mapsto w_{i}^{d_{i}},\text{ }1\leq i\leq s,\quad z_{i}\mapsto w_{i},\text{ }i>s.$
Then, the root stack $\mathscr{X}_{(X,D,\boldsymbol{d})}$ given in Theorem 4.3 is
$\mathscr{X}_{(X,D,\boldsymbol{d})}\cong[{\rm Spec}(k[w_{1},\dots,w_{d}])/\mu_{d_{1}}\times\cdots\times\mu_{d_{s}}],$
where $\mu_{d_{i}}$ is the group scheme of $d_{i}$-th roots of unity.
In the following, if there is no ambiguity, we use the notation $\mathscr{X}$ for the root stack $\mathscr{X}_{(X,D,\boldsymbol{d})}$ given in Theorem 4.3. With the same notation as above, let $\eta_{i}$ be the generic point of $D_{i}$. We define a discrete valuation ring $R_{i}:=\mathcal{O}_{X,\eta_{i}}$ with fraction field $K_{i}$. Denote by $\widehat{R}_{i}$ and $\widehat{K}_{i}$ the corresponding completions. More generally, if $R$ is a local ring, we always use the notation $\widehat{R}$ for the completion and $\widehat{K}$ for the fraction field. There exists a discrete valuation ring $\widetilde{R}_{i}$ such that
$\mathscr{X}\times_{X}{\rm Spec}\,R_{i}\cong[{\rm Spec}\,\widetilde{R}_{i}/\mu_{d_{i}}].$
Furthermore, let $w_{i}$ (resp. $z_{i}$) be a generator of the maximal ideal of $\widetilde{R}_{i}$ (resp. $R_{i}$), and then the isomorphism $\mathscr{X}\times_{X}{\rm Spec}\,R_{i}\cong[{\rm Spec}\,\widetilde{R}_{i}/\mu_{d_{i}}]$ is induced by a morphism $R_{i}\rightarrow\widetilde{R}_{i}$ such that $z_{i}$ is sent to $w_{i}^{d_{i}}$. Based on this property, we give the definition of type of $G$-bundles $\mathscr{E}$ on $\mathscr{X}$. The scheme $\mathscr{E}_{i}:=\mathscr{E}\times_{\mathscr{X}}[{\rm Spec}\,\widetilde{R}_{i}/\mu_{d_{i}}]$ can be regarded as a $\mu_{d_{i}}$-equivariant $G$-bundle on ${\rm Spec}\,\widetilde{R}_{i}$. Since $\widetilde{R}_{i}$ is a discrete valuation ring, a $\mu_{d_{i}}$-action on $\mathscr{E}_{i}$ is given by a representation $\rho_{i}:\mu_{d_{i}}\rightarrow T$ (see [BS15, §2.2.4] or [BP22, §5] for instance).
Definition 4.5.
Given a collection of representations $\boldsymbol{\rho}=\{\rho_{i}:\mu_{d_{i}}\rightarrow T,1\leq i\leq s\}$, a $G$-bundle $\mathscr{E}$ on $\mathscr{X}$ is of type $\boldsymbol{\rho}$ if the $\mu_{d_{i}}$-action on $\mathscr{E}_{i}$ is given by $\rho_{i}$.
In the following, we give an equivalent description of $\mathscr{E}$ via patching local data and the main technique we use comes from [BLR90, chapter 6]. It is natural to see that a $G$-bundle $\mathscr{E}$ gives a tuple $(\mathscr{E}_{0},\widehat{\mathscr{E}}_{i},\Theta_{i})$, where
$\mathscr{E}_{0}:=\mathscr{E}\times_{\mathscr{X}}\mathscr{X}\backslash\widetilde{D},\quad\widehat{\mathscr{E}}_{i}:=\mathscr{E}_{i}\times_{[{\rm Spec}\,\widetilde{R}_{i}/\mu_{d_{i}}]}[{\rm Spec}\,\widehat{\widetilde{R}}_{i}/\mu_{d_{i}}]$
and
$\Theta_{i}:\widehat{\mathscr{E}}_{i}\times_{[{\rm Spec}\,\widehat{\widetilde{R}}_{i}/\mu_{d_{i}}]}[{\rm Spec}\,\widehat{\widetilde{K}}_{i}/\mu_{d_{i}}]\xrightarrow{\cong}\mathscr{E}_{0}\times_{\mathscr{X}\backslash\widetilde{D}}[{\rm Spec}\,\widehat{\widetilde{K}}_{i}/\mu_{d_{i}}]$
is an isomorphism for each $i$. Equivalently, the isomorphism $\Theta_{i}$ can be regarded as an element in $G(\widehat{\widetilde{K}}_{i})$ such that
$\Theta_{i}(\gamma_{i}\cdot w_{i})=\rho_{i}(\gamma_{i})\Theta_{i}(w_{i}).$
Denote by $G(\widehat{\widetilde{K}}_{i})^{\rho_{i}}\subseteq G(\widehat{\widetilde{K}}_{i})$ the subset
$G(\widehat{\widetilde{K}}_{i})^{\rho_{i}}:=\{\Theta_{i}\in G(\widehat{\widetilde{K}}_{i})\,|\,\Theta_{i}(\gamma_{i}\cdot w_{i})=\rho_{i}(\gamma_{i})\Theta_{i}(w_{i})\}.$
Similarly, we define a subset $G(\widehat{\widetilde{R}}_{i})^{\rho_{i}}\subseteq G(\widehat{\widetilde{R}}_{i})$ as follows
$G(\widehat{\widetilde{R}}_{i})^{\rho_{i}}:=\{\Psi_{i}\in G(\widehat{\widetilde{R}}_{i})\,|\,\rho_{i}(\gamma_{i})\Psi_{i}(w_{i})\rho_{i}(\gamma_{i})^{-1}=\Psi_{i}(\gamma_{i}\cdot w_{i})\}.$
We consider the other direction that given a tuple $(\mathscr{E}_{0},\widehat{\mathscr{E}}_{i},\Theta_{i})$, where
  • $\mathscr{E}_{0}$ is a $G$-bundle on $\mathscr{X}\backslash\widetilde{D}$,
  • ^$\widehat{\mathscr{E}}_{i}$ is a $G$-bundle on $[{\rm Spec}\,\widehat{\widetilde{R}}_{i}/\mu_{d_{i}}]$,
  • $\Theta_{i}:\mathscr{E}_{0}\times_{\mathscr{X}\backslash\widetilde{D}}[{\rm Spec}\,\widehat{\widetilde{K}}_{i}/\mu_{d_{i}}]\rightarrow\widehat{\mathscr{E}}_{i}\times_{[{\rm Spec}\,\widehat{\widetilde{R}}_{i}/\mu_{d_{i}}]}[{\rm Spec}\,\widehat{\widetilde{K}}_{i}/\mu_{d_{i}}]$ is an isomorphism,
we want to construct a unique $G$-bundle (up to isomorphism) on $\mathscr{X}$. If $D_{i}$ are pairwise disjoint, then we can obtain a $G$-bundle on $\mathscr{X}$ by applying the argument of [BLR90, 6.2 Proposition D.4] for root stacks. In the general case, the data $(\mathscr{E}_{0},\widehat{\mathscr{E}}_{i},\Theta_{i})$ is not enough to give a $G$-bundle on $\mathscr{X}$ and the main problem comes from the intersections $D_{i}\cap D_{j}$. In order to obtain a $G$-bundle on $\mathscr{X}$, the isomorphisms $\{\Theta_{i},1\leq i\leq s\}$ should be “compatible” on the intersections. In the following, we give a precise description of the compatibility condition.
If the intersection of $D_{i}$ and $D_{j}$ is not empty, we define an algebraic cycle $D_{ij}=D_{i}\cdot D_{j}$ and denote by $\eta_{ij}$ its generic point. Let $R_{ij}:=\mathcal{O}_{X,\eta_{ij}}$ be the localization with fraction field $K_{ij}$. Similarly, there exists a ring $\widetilde{R}_{ij}$ such that $\mathscr{X}\times_{X}{\rm Spec}\,R_{ij}\cong[{\rm Spec}\,\widetilde{R}_{ij}/\mu_{d_{i}}\times\mu_{d_{j}}]$. For each pair $i,j$, if there is an isomorphism
$\Theta_{ij}:\widehat{\mathscr{E}_{i}}\times_{[{\rm Spec}\,\widehat{\widetilde{R}}_{ij}/\mu_{d_{i}}\times\mu_{d_{j}}]}[{\rm Spec}\,\widehat{\widetilde{K}}_{ij}/\mu_{d_{i}}\times\mu_{d_{j}}]\xrightarrow{\cong}\widehat{\mathscr{E}_{j}}\times_{[{\rm Spec}\,\widehat{\widetilde{R}}_{ij}/\mu_{d_{i}}\times\mu_{d_{j}}]}[{\rm Spec}\,\widehat{\widetilde{K}}_{ij}/\mu_{d_{i}}\times\mu_{d_{j}}]$
such that $(p_{0i}\circ\Theta_{i})(s_{i})=(p_{0j}\circ\Theta_{j})(s_{j})$ if and only if $(\Theta_{ij}\circ p_{ij})(s_{i})=p_{ji}(s_{j})$ for any section $s_{i}$ (resp. $s_{j}$) of ^$\widehat{\mathscr{E}}_{i}\times_{[{\rm Spec}\,\widehat{\widetilde{R}}_{i}/\mu_{d_{i}}]}[{\rm Spec}\,\widehat{\widetilde{K}}_{i}/\mu_{d_{i}}]$ (resp. ^$\widehat{\mathscr{E}}_{j}\times_{[{\rm Spec}\,\widehat{\widetilde{R}}_{j}/\mu_{d_{j}}]}[{\rm Spec}\,\widehat{\widetilde{K}}_{j}/\mu_{d_{j}}]$), then we say that the tuple (,^$(\mathscr{E},\widehat{\mathscr{E}}_{i},\Theta_{i})$ is compatible.
${\mathscr{E}_{0}\times_{\mathscr{X}\backslash\widetilde{D}}[{\rm Spec}\,\widehat{\widetilde{K}}_{ij}/\mu_{d_{i}}\times\mu_{d_{j}}]}$${\mathscr{E}_{0}\times_{\mathscr{X}\backslash\widetilde{D}}[{\rm Spec}\,\widehat{\widetilde{K}}_{i}/\mu_{d_{i}}]}$${\mathscr{E}_{0}\times_{\mathscr{X}\backslash\widetilde{D}}[{\rm Spec}\,\widehat{\widetilde{K}}_{j}/\mu_{d_{j}}]}$^${\widehat{\mathscr{E}}_{i}\times_{[{\rm Spec}\,\widehat{\widetilde{R}}_{i}/\mu_{d_{i}}]}[{\rm Spec}\,\widehat{\widetilde{K}}_{i}/\mu_{d_{i}}]}$^${\widehat{\mathscr{E}}_{j}\times_{[{\rm Spec}\,\widehat{\widetilde{R}}_{j}/\mu_{d_{j}}]}[{\rm Spec}\,\widehat{\widetilde{K}}_{j}/\mu_{d_{j}}]}$^${\widehat{\mathscr{E}}_{i}\times_{[{\rm Spec}\,\widehat{\widetilde{R}}_{ij}/\mu_{d_{i}}\times\mu_{d_{j}}]}[{\rm Spec}\,\widehat{\widetilde{K}}_{ij}/\mu_{d_{i}}\times\mu_{d_{j}}]}$^${\widehat{\mathscr{E}}_{j}\times_{[{\rm Spec}\,\widehat{\widetilde{R}}_{ij}/\mu_{d_{i}}\times\mu_{d_{j}}]}[{\rm Spec}\,\widehat{\widetilde{K}}_{ij}/\mu_{d_{i}}\times\mu_{d_{j}}]}$$\scriptstyle{p_{0i}}$$\scriptstyle{p_{0j}}$$\scriptstyle{\Theta_{i}}$$\scriptstyle{p_{ij}}$$\scriptstyle{\Theta_{j}}$$\scriptstyle{p_{ji}}$$\scriptstyle{\Theta_{ij}}$
It is routine to check that this will give a $G$-bundle on $\mathscr{X}$. The above discussion gives the following lemma.
Lemma 4.6.
A $G$-bundle $\mathscr{E}$ on $\mathscr{X}$ is equivalent to a compatible tuple (,~$(\mathscr{E},\widetilde{\mathscr{E}}_{i},\Theta_{i})$.
Remark 4.7.
In the case of curves, $\widetilde{D}$ is a set of distinct points. Thus, the intersection $\widetilde{D}_{i}\cap\widetilde{D}_{j}$ is always empty. In this case, isomorphisms
$\Theta_{i}:\widehat{\mathscr{E}}_{i}\times_{[{\rm Spec}\,\widehat{\widetilde{R}}_{i}/\mu_{d_{i}}]}[{\rm Spec}\,\widehat{\widetilde{K}}_{i}/\mu_{d_{i}}]\xrightarrow{\cong}\mathscr{E}_{0}\times_{\mathscr{X}\backslash\widetilde{D}}[{\rm Spec}\,\widehat{\widetilde{K}}_{i}/\mu_{d_{i}}]$
give all information of a $G$-bundle $\mathscr{E}$. Moreover, isomorphism classes of $G$-bundles on $\mathscr{X}$ of type $\boldsymbol{\rho}$ are in one-to-one correspondence with the double coset
$\left[G(k[X\backslash D])\backslash\prod_{i=1}^{s}G(\widehat{\widetilde{K}}_{i})^{\rho_{d_{i}}}/\prod_{i=1}^{s}G(\widehat{\widetilde{R}}_{i})^{\rho_{d_{i}}}\right],$
and this is an adelic description of $G$-bundles in the case of curves [BS15, §3].
4.3. Logarithmic G-Higgs Bundles and Logarithmic Flat G-bundles
In the previous subsection, a $G$-bundle on $\mathscr{X}\cong[Y/R]$ is regarded as one on $Y$ equipped with an isomorphism $\tau$. In this subsection, we define logarithmic $G$-Higgs bundles and logarithmic flat $G$-bundles on $\mathscr{X}=\mathscr{X}_{(X,D,\boldsymbol{d})}$ in the same way as what we defined on varieties in §2.1.
Definition 4.8.
A logarithmic $G$-Higgs bundle on $\mathscr{X}$ is a pair $(\mathscr{E},\theta)$, where $\mathscr{E}$ is a $G$-bundle on $\mathscr{X}$ and $\theta:\mathscr{X}\rightarrow\mathscr{E}(\mathfrak{g})\otimes\Omega_{\mathscr{X}}({\rm log}\,\widetilde{D})$ is integrable. A logarithmic flat $G$-bundle on $\mathscr{X}$ is a pair $(\mathscr{V},\nabla)$, where $\mathscr{V}$ is a $G$-bundle $\mathscr{X}$ and $\nabla:\mathcal{O}_{\mathscr{V}}\rightarrow\mathcal{O}_{\mathscr{V}}\otimes\Omega_{\mathscr{X}}({\rm log}\,\widetilde{D})$ is an integrable connection.
Let ${\rm HIG}_{p-1}(\mathscr{X}_{\rm log}/k,G)$ be the category of nilpotent logarithmic $G$-Higgs bundles on $\mathscr{X}$ of exponent $\leq p-1$ on $\mathscr{X}$. For nilpotent logarithmic flat $G$-bundles, note first that we have a short exact sequence
$0\longrightarrow\Omega_{\mathscr{X}}\longrightarrow\Omega_{\mathscr{X}}({\rm log}\,\widetilde{D})\xrightarrow{\oplus{\rm res}_{\widetilde{D}_{i}}}\bigoplus_{i}\mathcal{O}_{\widetilde{D}_{i}}\longrightarrow 0$
Given a logarithmic flat $G$-bundle $(\mathscr{V},\nabla)$, we define
${\rm Res}_{\widetilde{D}_{i}}\nabla:\mathcal{O}_{\mathscr{V}}\otimes\mathcal{O}_{\widetilde{D}_{i}}\rightarrow\mathcal{O}_{\mathscr{V}}\otimes\mathcal{O}_{\widetilde{D}_{i}}$
as the composite
$\mathcal{O}_{\mathscr{V}}\xrightarrow{\nabla}\mathcal{O}_{\mathscr{V}}\otimes\Omega_{\mathscr{X}}({\rm log}\widetilde{D})\xrightarrow{{\rm id}\otimes{\rm res}_{\widetilde{D}_{i}}}\mathcal{O}_{\mathscr{V}}\otimes\mathcal{O}_{\widetilde{D}_{i}}.$
Let ${\rm MIC}_{p-1}(\mathscr{X}_{\rm log}/k,G)$ be the category of nilpotent logarithmic flat $G$-bundles of exponent $\leq p-1$ on $\mathscr{X}$ such that ${\rm Res}_{\widetilde{D}_{i}}\nabla$ is also nilpotent of exponent $\leq p-1$ for every $1\leq i\leq s$.
Remark 4.9.
Since the objects we consider here are $G$-bundles, the “tor condition” considered in [LSYZ19, Appendix] is satisfied automatically, i.e.
$\mathcal{T}or_{1}(\mathcal{O}_{\mathscr{E}},(F_{1})_{*}\mathcal{O}_{\widetilde{D}_{i}})=0,\,\mathcal{T}or_{1}(\mathcal{O}_{\mathscr{V}},\mathcal{O}_{\widetilde{D}_{i}})=0,$
where $\mathscr{E}$ is a $G$-bundle on $\mathscr{X}^{\prime}$ and $\mathscr{V}$ is a $G$-bundle on $\mathscr{X}$.
Lemma 4.10.
Suppose that $\mathscr{X}$ is $W_{2}(k)$-liftable. The following categories are equivalent
${{\rm HIG}_{0}(\mathscr{X}^{\prime}_{\rm log}/k,G)}$${{\rm MIC}_{0}(\mathscr{X}_{\rm log}/k,G)}$${{\rm HIG}_{0}(\mathscr{X}^{\prime}/k,G)}$${{\rm MIC}_{0}(\mathscr{X}/k,G)\,,}$
where we use ``="``"``="` ` = " for equivalence.
Proof.
The proof is a stacky version of [LSYZ19, Lemma A.2]. For the Higgs side, the equivalence
${\rm HIG}_{0}(\mathscr{X}^{\prime}_{\rm log}/k,G)={\rm HIG}_{0}(\mathscr{X}^{\prime}/k,G)$
is clear because Higgs fields are trivial in this case. Now we take an object $(\mathscr{V},\nabla)\in{\rm MIC}_{0}(\mathscr{X}_{\rm log}/k,G)$. Since $\mathcal{T}or_{1}(\mathcal{O}_{\mathscr{V}},\mathcal{O}_{\widetilde{D}_{i}})=0$, we have a short exact sequence
$0\rightarrow\mathcal{O}_{\mathscr{V}}\otimes\Omega_{\mathscr{X}}\rightarrow\mathcal{O}_{\mathscr{V}}\otimes\Omega_{\mathscr{X}}({\rm log}\,\widetilde{D})\xrightarrow{{\rm id}\otimes{\rm res}}\bigoplus_{i}\mathcal{O}_{\mathscr{V}}\otimes\mathcal{O}_{\widetilde{D}_{i}}\longrightarrow 0$
Since the residue is nilpotent of exponent $\leq 0$, it is trivial and we have $\nabla(\mathcal{O}_{\mathscr{V}})\subseteq\mathcal{O}_{\mathscr{V}}\otimes\Omega_{\mathscr{X}}$. Therefore,
${\rm MIC}_{0}(\mathscr{X}_{\rm log}/k,G)={\rm MIC}_{0}(\mathscr{X}/k,G).$
The equivalence
${\rm HIG}_{0}(\mathscr{X}^{\prime}/k,G)={\rm MIC}_{0}(\mathscr{X}/k,G)$
follows from Proposition 4.1 directly. This finishes the proof of this lemma. ∎
Lemma 4.11.
The following diagram commutes.
[…]
Proof.
To prove this lemma, it is enough to work around a neighborhood of $\widetilde{x}\in\widetilde{D}$. By Theorem 4.3 and Example 4.4, we can choose a neighborhood around $\widetilde{x}$ such that it is isomorphic to a root stack in the form $\mathscr{X}_{(X,D,\boldsymbol{d})}\cong[{\rm Spec}(k[w_{1},\dots,w_{d}])/\mu_{d_{1}}\times\cdots\times\mu_{d_{s}}]$. Then, this lemma follows directly from [LSYZ19, Claim A.3]. ∎
Theorem 4.12.
Suppose that $(\mathscr{X},\widetilde{D})$ is $W_{2}$-liftable. Then the categories ${\rm HIG}_{p-1}(\mathscr{X}^{\prime}_{\rm log}/k,G)$ and ${\rm MIC}_{p-1}(\mathscr{X}_{\rm log}/k,G)$ are equivalent.
Proof.
Similar to the proof of [LSYZ19, Theorem A.1], based on Theorem 4.2 and Lemma 4.10, it is enough to work on local charts, and thus we suppose that
$\mathscr{X}=[{\rm Spec}(k[w_{1},\dots,w_{d}])/\mu_{d_{1}}\times\cdots\times\mu_{d_{s}}].$
Given $(\mathscr{E},\theta)\in{\rm HIG}_{p-1}(\mathscr{X}^{\prime}_{\rm log}/k,G)$, we obtain a logarithmic flat $G$-bundle $(\mathscr{V},\nabla)$ via a logarithmic version of the inverse Cartier transform given in Theorem 4.2. For convenience, we regard $(\mathscr{E},\theta)$ (resp. $(\mathscr{V},\nabla)$) as equivariant logarithmic $G$-Higgs bundles $(E,\theta)$ (resp. logarithmic flat $G$-bundles $(V,\nabla)$) on ${\rm Spec}(k[y_{1},\dots,y_{d}])$. Thus, following the construction of (3.2) in the proof of Theorem 3.1, we have
$\nabla=\nabla_{can}+\zeta(F_{1}^{*}\theta),$
and then
${\rm Res}_{\widetilde{D}_{i}}\nabla=F_{1}^{*}{\rm Res}_{\widetilde{D}_{i}}\theta.$
This implies that the residues are nilpotent. Therefore, we have $(\mathscr{V},\nabla)\in{\rm MIC}_{p-1}(\mathscr{X}_{\rm log}/k,G)$.
For the other direction, we take an object $(\mathscr{V},\nabla)\in{\rm MIC}_{p-1}(\mathscr{X}_{\rm log}/k,G)$. With the same proof as in §3.2, we obtain a triple $(\mathscr{V}^{\prime},\nabla^{\prime},\psi^{\prime})$, where $(\mathscr{V}^{\prime},\nabla^{\prime})$ is a logarithmic flat $G$-bundle with trivial $p$-curvature and $\psi^{\prime}$ is the $F$-Higgs field horizontal with respect to $\nabla^{\prime}$. Then, the triple $(\mathscr{V}^{\prime},\nabla^{\prime},\psi^{\prime})$ descents to a nilpotent logarithmic $G$-Higgs bundle $(\mathscr{E},\theta)$ on $\mathscr{X}^{\prime}$, which is clear an object in the category ${\rm HIG}_{p-1}(\mathscr{X}^{\prime}_{\rm log}/k,G)$. ∎
Remark 4.13.
In this remark, we discuss how type changes via the correspondence given in Theorem 4.12. Since type is a local property, we still work on the stack
$\mathscr{X}=[{\rm Spec}(k[w_{1},\dots,w_{d}])/\mu_{d_{1}}\times\cdots\times\mu_{d_{s}}]$
as in Example 4.4 and consider equivariant objects. We fix some notations in this example. Let
$\Gamma=\mu_{d_{1}}\times\cdots\times\mu_{d_{s}},\quad Y={\rm Spec}(k[w_{1},\dots,w_{d}]).$
Let $E\cong Y^{\prime}\times G$ be a trivial $G$-bundle of type $\boldsymbol{\rho}=\{\rho_{i},1\leq i\leq s\}$ on $Y$. Thus, the $\Gamma$-action on $E$ is given as follows
$(\gamma_{1},\dots,\gamma_{s})\cdot((w_{1},\dots,w_{d}),g)=((\gamma_{1}w_{1},\dots,\gamma_{s}w_{s},\dots,w_{d}),\rho_{1}(\gamma_{1})\dots\rho_{s}(\gamma_{s})g).$
[…]
4.4. Stability Condition
In this subsection, we make a further assumption on the root stack $\mathscr{X}$ that it is projective. In other words, the coarse moduli space $X$ is projective. We fix an ample line bundle $H$ on $X$ and a locally free sheaf $\mathscr{L}$ on $\mathscr{X}$. The modified Hilbert polynomial $P_{\mathscr{L}}(\mathscr{F},m)$ is defined as follows:
$P_{\mathscr{L}}(\mathscr{F},m)=\chi(\mathscr{X},\mathscr{F}\otimes\mathscr{L}\otimes\pi^{*}(H^{m})),\,m\gg 0.$
When $\mathscr{F}$ is pure of dimension $d$, the function $P_{\mathscr{L}}(\mathscr{F},m)$ can be written as
$P_{\mathscr{L}}(\mathscr{F},m)=\sum_{i=0}^{d}\alpha_{\mathscr{L},i}(\mathscr{F})\frac{m^{i}}{i!},$
and the $\mathscr{L}$-degree of $\mathscr{F}$ is given as
$\frac{\deg_{\mathscr{L}}\mathscr{F}}{r}=\frac{\alpha_{\mathscr{L},d-1}(\mathscr{F})}{re}-\frac{\alpha_{\mathscr{L},d-1}(\mathcal{O}_{\mathscr{X}})}{e},$
where $r$ is the rank of $\mathscr{F}$ and $e$ is the rank of $\mathscr{L}$ [Nir09, Proposition 3.18]. As pointed out in [Nir09, Corollary 7.10], with a good choice of $\mathscr{L}$, the degree $\deg_{\mathscr{L}}\mathscr{F}$ is equal to the parabolic degree of the corresponding parabolic sheaf on $X$. We also refer the reader to [Bor07] for more details.
Remark 4.14.
If $\mathscr{X}$ is a global quotient, the locally free sheaf $\mathscr{L}$ is a generating sheaf [Kre09, Theorem 5.3]. Moreover, generating sheaves are very helpful to construct the moduli spaces of coherent sheaves and Higgs sheaves on Deligne–Mumford stacks (see [OS03, Nir09, Sun20, Sun22] for instance).
We define stability conditions for logarithmic $G$-Higgs bundles and logarithmic flat $G$-bundles on $\mathscr{X}$ similar as those in §2.3.
Definition 4.15.
A logarithmic $G$-Higgs bundle $(\mathscr{E},\theta)$ is $R$-semistable (resp. $R$-stable), if
  • for any proper parabolic subgroup $P\subseteq G$;
  • for any antidominant character $\chi:P\rightarrow\mathbb{G}_{m}$ acting trivially on the center;
  • for any $\theta$-compatible reduction of structure group $\sigma:\mathscr{X}\rightarrow\mathscr{E}/P$,
we have
[…]
Denote by ${\rm HIG}^{s(s)}_{p-1}(\mathscr{X}_{\rm log}/k,G)\subseteq{\rm HIG}_{p-1}(\mathscr{X}^{\prime}_{\rm log}/k,G)$ the subcategory of $R$-(semi)stable nilpotent logarithmic $G$-Higgs bundles of exponent $\leq p-1$ on $\mathscr{X}$.
A logarithmic flat $G$-bundle $(\mathscr{V},\nabla)$ is $R$-semistable (resp. $R$-stable), if
  • for any pparabolic subgroup $P\subseteq G$;
  • for any antidominant character $\chi:P\rightarrow\mathbb{G}_{m}$ acting trivially on the center;
  • for any $\nabla$-compatible reduction of structure group $\varsigma:\mathscr{X}\rightarrow\mathscr{V}/P$,
we have
[…]
Let ${\rm MIC}^{(s)s}_{p-1}(\mathscr{X}_{\rm log}/k,G)$ be the subcategory of ${\rm MIC}_{p-1}(\mathscr{X}_{\rm log}/k,G)$ consisting of $R$-(semi)stable nilpotent logarithmic flat $G$-bundles of exponent $\leq p-1$ on $\mathscr{X}$.
Similar to the setup in §3.3, we define $\mathscr{L}^{\prime}:=\pi^{*}_{\mathscr{X}/k}\mathscr{L}$ and $H^{\prime}:=\pi^{*}_{X/k}H$. Then the degree of a pure sheaf $\mathcal{F}^{\prime}$ on $\mathscr{X}^{\prime}$ is given with respect to $H^{\prime}$ and $\mathscr{F}^{\prime}$. It is easy to check that $\deg_{\mathscr{L}^{\prime}}\mathscr{F}^{\prime}=p\deg_{\mathscr{L}}\mathscr{F}$. Then, the stability condition of Higgs bundles on $\mathscr{X}^{\prime}$ is defined by $H^{\prime}$ and $\mathscr{F}^{\prime}$, while the stability condition of flat bundles on $\mathscr{X}$ is given with respect to $H$ and $\mathscr{F}$.
Theorem 4.16.
Suppose that $(\mathscr{X},\widetilde{D})$ is $W_{2}(k)$-liftable. The categories ${\rm HIG}^{(s)s}_{p-1}(\mathscr{X}^{\prime}_{\rm log}/k,G)$ and ${\rm MIC}^{(s)s}_{p-1}(\mathscr{X}_{\rm log}/k,G)$ are equivalent.
Proof.
Given $(\mathscr{E},\theta)\in{\rm HIG}_{p-1}(\mathscr{X}^{\prime}_{\rm log}/k,G)$, Theorem 4.12 shows that there is a logarithmic flat $G$-bundle $(\mathscr{V},\nabla)$ in ${\rm MIC}_{p-1}(\mathscr{X}_{\rm log}/k,G)$. Then, the proof is similar to that of Theorem 3.6 and we only have to check that there is a one-to-one correspondence between $\theta$-compatible reductions of $\mathscr{E}$ and $\nabla$-compatible reductions of $\mathscr{V}$. In this proof, we start with a $\theta$-compatible reduction $\sigma$ and construct a $\nabla$-compatible reduction $\varsigma$. The other direction can be proved in a similar way and we omit the details.
Let $(\mathscr{E},\theta)\in{\rm HIG}_{p-1}(\mathscr{X}^{\prime}_{\rm log}/k,G)$ be a $G$-Higgs bundle. Under the isomorphism $\mathscr{X}\cong[Y/R]$, we regard $(\mathscr{E},\theta)$ as a tuple $(E,\theta,\tau)$, where $(E,\theta)$ is a logarithmic $G$-Higgs bundle on $Y^{\prime}$ and $\tau:p_{1}^{*}(E,\theta)\cong p_{2}^{*}(E,\theta)$ is an isomorphism as in the proof of Theorem 4.2. A $\theta$-compatible reduction $\sigma:\mathscr{X}^{\prime}\rightarrow\mathscr{E}/P$ can be regarded as a $\theta$-compatible reduction $\sigma:Y^{\prime}\rightarrow E/P$ such that the following diagram commutes,
${p_{1}^{*}E_{\sigma}}$${p_{1}^{*}E}$${p_{2}^{*}E_{\sigma}}$${p_{2}^{*}E}$$\scriptstyle{p_{1}^{*}\sigma}$$\scriptstyle{\tau_{\sigma}}$$\scriptstyle{\tau}$$\scriptstyle{p_{2}^{*}\sigma}$
where $\tau_{\sigma}:p_{1}^{*}E_{\sigma}\rightarrow p_{2}^{*}E_{\sigma}$ is the isomorphism induced by $\tau$. Moreover, $p_{i}^{*}\sigma$ is compatible with the Higgs field $p_{i}^{*}\theta$. By Theorem 3.6, the triple $(E,\theta,\sigma)$ on $Y^{\prime}$ corresponds to a triple $(V,\nabla,\varsigma)$, where $(V,\nabla)$ is a nilpotent logarithmic flat $G$-bundle and $\varsigma:Y\rightarrow V/P$ is a $\nabla$-compatible reduction. Similarly, $p_{i}^{*}(E,\theta,\sigma)$ corresponds to $p_{i}^{*}(V,\nabla,\varsigma)$ on $R$. The isomorphisms
$\tau:p_{1}^{*}(E,\theta)\cong p_{2}^{*}(E,\theta),\quad\tau_{\sigma}:p_{1}^{*}(E_{\sigma},\theta_{\sigma})\cong p_{2}^{*}(E_{\sigma},\theta_{\sigma})$
induce ones
$F_{1}^{*}\tau:p_{1}^{*}(V,\nabla)\cong p_{2}^{*}(V,\nabla),\quad F_{1}^{*}\tau_{\sigma}:p_{1}^{*}(V_{\varsigma},\nabla_{\varsigma})\cong p_{2}^{*}(V_{\varsigma},\nabla_{\varsigma})$
by Theorem 4.2. Thus, we obtain a triple $(V,\nabla,\varsigma)$ together with an isomorphism $F_{1}^{*}\tau$ such that the diagram
[…]
commutes, and then we get a nilpotent logarithmic flat $G$-bundle $(\mathscr{V},\nabla)\in{\rm MIC}_{p-1}(\mathscr{X}_{\rm log}/k,G)$ together with a $\nabla$-compatible reduction $\varsigma$ of structure group. ∎
5. Logahoric Higgs Torsors and Logahoric Connections
The goal of this section is to prove a logahoric version of nonabelian Hodge correspondence for $G$-bundles via exponential twisting. In §5.1, we review the construction of parahoric torsors over a smooth variety $X$. The case on curves is studied by Balaji–Seshadri [BS15] and the higher dimensional case is considered by Balaji–Pandey recently [BP22]. In §5.2, we prove a correspondence between parahoric torsors on $X$ and $G$-bundles on a root stack $\mathscr{X}$ (Proposition 5.5), and the correspondence also preserves the stability condition (Proposition 5.9). In §5.3, we give the proof of the main result (Theorem 5.17).
5.1. Parahoric Group Scheme
Recall that $G$ is a split connected reductive linear algebraic group over $k$. We fix a maximal torus $T\subseteq G$ and denote by $\mathcal{R}$ the root system. Define $X(T):={\rm Hom}(T,\mathbb{G}_{m})$ and $Y(T):={\rm Hom}(\mathbb{G}_{m},T)$. Denote by
$\langle\cdot,\cdot\rangle:Y(T)\times X(T)\rightarrow\mathbb{Z}$
the canonical pairing, which can be naturally extended to $\mathbb{Q}$. A weight $\alpha$ is an element in $Y(T)\otimes_{\mathbb{Z}}\mathbb{Q}$, and let $d$ be the least common multiple of the denominators. Clearly, $d\alpha\in Y(T)$. For convenience, we say that $d$ is the denominator of $\alpha$. The weight $\alpha$ is called tame if $d$ is coprime to $p$ [LS24, §2]. Given a weight $\alpha$ and a root $r\in\mathcal{R}$, the pairing $\langle\alpha,r\rangle$ is a rational number, and sometimes we use the notation $r(\alpha)$. All weights considered in this paper are assumed to be tame and satisfy the condition $r(\alpha)\leq 1$ for any $r\in\mathcal{R}$.
Let $X$ be a smooth variety over $k$ with a given reduced normal crossing divisor $D=\sum\limits_{i=1}^{s}D_{i}$, where $D_{i}$ are irreducible components for $1\leq i\leq s$. We also equip each $D_{i}$ with a tame weight $\alpha_{i}$, and denote by $\boldsymbol{\alpha}$ the collection of these weights. Furthermore, denote by $\boldsymbol{d}=\{d_{i},1\leq i\leq s\}$ the collection of the corresponding denominators. Recall that $\eta_{i}$ is the generic point of $D_{i}$ and $R_{i}=\mathcal{O}_{X,\eta_{i}}$ with fraction field $K_{i}$. Denote by $\widehat{R}_{i}$ and (resp. $\widehat{K}_{i}$) the completion, which is regarded as the formal neighbourhood of the generic point $\eta_{i}$. Let $G_{\alpha_{i}}(\widehat{R}_{i})$ be the parahoric group associated to a given weight $\alpha_{i}$
$G_{\alpha_{i}}(\widehat{R}_{i}):=\langle T(\widehat{R}_{i}),U_{r}(z_{i}^{m_{r}(\alpha_{i})}\widehat{R}_{i}),r\in\mathcal{R}\rangle,$
[…]
Theorem 5.1 (Theorem 5.4 in [BP22]).
There exists a smooth affine group scheme $\mathcal{G}_{\boldsymbol{\alpha}}$ on $X$ such that
  • $\mathcal{G}_{\boldsymbol{\alpha}}|_{X\backslash D}\cong G\times X\backslash D$,
  • $\mathcal{G}_{\boldsymbol{\alpha}}|_{{\rm Spec}\,\widehat{R}_{i}}\cong\mathcal{G}_{\alpha_{i}}$.
This group scheme is unique up to isomorphism, and is called a parahoric group scheme. Moreover, there exists a vector bundle $\mathfrak{g}_{\boldsymbol{\alpha}}$ with a natural Lie algebra structure on its fiber such that
  • $\mathfrak{g}_{\boldsymbol{\alpha}}|_{X\backslash D}\cong\mathfrak{g}\times X\backslash D$,
  • $\mathfrak{g}_{\boldsymbol{\alpha}}|_{{\rm Spec}\,\widehat{R}_{i}}\cong\mathfrak{g}_{\alpha_{i}}$.
Remark 5.2.
There is a well-defined $\mathcal{G}_{\boldsymbol{\alpha}}$-action on $\mathfrak{g}_{\boldsymbol{\alpha}}$ induced by the adjoint action locally, and we refer the reader to [Boa11] for more details. Therefore, given a parahoric $\mathcal{G}_{\boldsymbol{\alpha}}$-torsor $\mathcal{E}$, we can define its adjoint bundle $\mathcal{E}(\mathfrak{g})$ naturally.
Example 5.3.
In this example, we give a more precise description about $\mathcal{G}_{\boldsymbol{\alpha}}$ around a formal neighborhood of a point $x\in D$. For each root $r\in\mathcal{R}$, we have a natural morphim
$u_{r}:\mathbb{G}_{a}\rightarrow G,$
and denote by $U_{r}$ the image, which is a unipotent group. Let $\widehat{\mathcal{O}}_{X,x}\cong k[[z_{1},\dots,z_{d}]]$ be a formal neighborhood of $x\in D\subseteq X$ with $\widehat{K}$ the fraction field. Suppose that $D_{i}=Z(z_{i})$ for $1\leq i\leq s$, and we assume that $s\leq d$ for convenience. Then, $\mathcal{G}_{\boldsymbol{\alpha}}|_{\widehat{\mathcal{O}}_{X,x}}$ is isomorphic to the group scheme on ${\rm Spec}\,\widehat{\mathcal{O}}_{X,x}$ determined by the following generators
$\langle T(\widehat{\mathcal{O}}),\,U_{r}((\prod_{i=1}^{s}z_{i}^{m_{r}(\alpha_{i})})\widehat{\mathcal{O}}),\,r\in\mathcal{R}\rangle.$
This group scheme $\mathcal{G}_{\boldsymbol{\alpha}}|_{\widehat{\mathcal{O}}_{X,x}}$ is called a n-parahoric group scheme [BP22, §5]. Similarly, $\mathfrak{g}_{\boldsymbol{\alpha}}|_{\widehat{\mathcal{O}}_{X,x}}$ is isomorphic to a Lie subalgebra in $\mathfrak{g}(\widehat{K})$ generated by
$\langle\mathfrak{t}(\widehat{\mathcal{O}}),\,\mathfrak{g}_{r}((\prod_{i=1}^{s}z_{i}^{m_{r}(\alpha_{i})})\widehat{\mathcal{O}}),\,r\in\mathcal{R}\rangle.$
5.2. Parahoric Torsors on $X$ vs. Principal bundles on $\mathscr{X}$
We fix the following data first:
  • $X$: smooth variety,
  • $D$: reduced normal crossing divisor with irreducible components $D_{i}$, $1\leq i\leq s$,
  • $\boldsymbol{\alpha}$: a collection of tame weights.
By Theorem 5.1, we obtain a parahoric group scheme $\mathcal{G}_{\boldsymbol{\alpha}}$ on $X$. Now we consider parahoric $\mathcal{G}_{\boldsymbol{\alpha}}$-torsors. Given a parahoric $\mathcal{G}_{\boldsymbol{\alpha}}$-torsor, we associate it with a tuple $(\mathcal{E}_{0},\widehat{\mathcal{E}}_{i},\Theta_{i})$, where
$\mathcal{E}_{0}:=\mathcal{E}\times_{X}X\backslash D,\quad\widehat{\mathcal{E}}_{i}:=\mathcal{E}\times_{X}{\rm Spec}\,\widehat{R}_{i}$
and
$\Theta_{i}:\widehat{\mathcal{E}}_{i}\times_{{\rm Spec}\,\widehat{R}_{i}}{\rm Spec}\,\widehat{K}_{i}\rightarrow\mathcal{E}_{0}\times_{X\backslash D}{\rm Spec}\,\widehat{K}_{i},$
is an isomorphism for each $i$. In this case, the isomorphism $\Theta_{i}$ can be regarded as an element in $G(\widehat{K}_{i})$.
For the other direction, we start with a such a tuple $(\mathcal{E}_{0},\widehat{\mathcal{E}}_{i},\Theta_{i})$, where
  • $\mathcal{E}_{0}$ is a $G$-bundle on $X\backslash D$;
  • ^$\widehat{\mathcal{E}}_{i}$ is a parahoric $\mathcal{G}_{\alpha_{i}}$-torsor on ${\rm Spec}\,\widehat{R}_{i}$;
  • $\Theta_{i}:\widehat{\mathcal{E}}_{i}\times_{{\rm Spec}\,\widehat{R}_{i}}{\rm Spec}\,\widehat{K}_{i}\rightarrow\mathcal{E}_{0}\times_{X\backslash D}{\rm Spec}\,\widehat{K}_{i}$ is an isomorphism,
and we want to construct a parahoric $\mathcal{G}_{\boldsymbol{\alpha}}$-torsor on $X$. Similar to the discussion in §4.2, if the intersection $D_{i}\cap D_{j}$ is nonempty for some distinct $i,j$, such a tuple $(\mathcal{E}_{0},\widehat{\mathcal{E}}_{i},\Theta_{i})$ may not determine a parahoric $\mathcal{G}_{\boldsymbol{\alpha}}$-torsor. We follow the same notations as in §4.2 and introduce the compatiblility condition in a similar way. Given such a tuple $(\mathcal{E}_{0},\widehat{\mathcal{E}}_{i},\Theta_{i})$, if there is an isomorphism
$\Theta_{ij}:\widehat{\mathcal{E}}_{i}\times_{{\rm Spec}\,\widehat{R}_{ij}}{\rm Spec}\,\widehat{K}_{ij}\rightarrow\widehat{\mathcal{E}}_{j}\times_{{\rm Spec}\,\widehat{R}_{ij}}{\rm Spec}\,\widehat{K}_{ij}$
such that $(q_{0i}\circ\Theta_{i})(s_{i})=(q_{0j}\circ\Theta_{j})(s_{j})$ if and only if $(\Theta_{ij}\circ q_{ij})(s_{i})=q_{ji}(s_{j})$ for any section $s_{i}$ (resp. $s_{j}$) of ^$\widehat{\mathcal{E}}_{i}\times_{{\rm Spec}\,\widehat{R}_{i}}{\rm Spec}\,\widehat{K}_{i}$ (resp. ^$\widehat{\mathcal{E}}_{j}\times_{{\rm Spec}\,\widehat{R}_{j}}{\rm Spec}\,\widehat{K}_{j}$), then we say that the tuple (,^$(\mathcal{E},\widehat{\mathcal{E}}_{i},\Theta_{i})$ is compatible.
${\mathcal{E}_{0}\times_{X\backslash D}{\rm Spec}\,\widehat{K}_{ij}}$${\mathcal{E}_{0}\times_{X\backslash D}{\rm Spec}\,\widehat{K}_{i}}$${\mathcal{E}_{0}\times_{X\backslash D}{\rm Spec}\,\widehat{K}_{j}}$^${\widehat{\mathcal{E}}_{i}\times_{{\rm Spec}\,\widehat{R}_{i}}{\rm Spec}\,\widehat{K}_{i}}$^${\widehat{\mathcal{E}}_{j}\times_{{\rm Spec}\,\widehat{R}_{j}}{\rm Spec}\,\widehat{K}_{j}}$^${\widehat{\mathcal{E}}_{i}\times_{{\rm Spec}\,\widehat{R}_{ij}}{\rm Spec}\,\widehat{K}_{ij}}$^${\widehat{\mathcal{E}}_{j}\times_{{\rm Spec}\,\widehat{R}_{ij}}{\rm Spec}\,\widehat{K}_{ij}}$$\scriptstyle{q_{0i}}$$\scriptstyle{p_{0j}}$$\scriptstyle{q_{ij}}$$\scriptstyle{\Theta_{i}}$$\scriptstyle{q_{ji}}$$\scriptstyle{\Theta_{j}}$$\scriptstyle{\Theta_{ij}}$
As an analogue of Lemma 4.6, we have the following lemma for parahoric torsors.
Lemma 5.4.
A parahoric $\mathcal{G}_{\boldsymbol{\alpha}}$-torsor $\mathcal{E}$ is equivalent to a compatible tuple (,~$(\mathcal{E},\widetilde{\mathcal{E}}_{i},\Theta_{i})$.
Now we move to principal bundles on root stacks. Let $\alpha$ be a weight and we suppose that its denominator is $d$. This weight $\alpha$ uniquely determines a representation $\rho:\mu_{d}\rightarrow T$ [BS15, Lemma 2.2.8]. Given a collection of weights $\boldsymbol{\alpha}$ with $\boldsymbol{d}$ the corresponding collection of denominators, they determine a collection of representations $\boldsymbol{\rho}$. Since all weights in $\boldsymbol{\alpha}$ are tame, we obtain a (tame) root stack $\mathscr{X}:=\mathscr{X}_{(X,D,\boldsymbol{d})}$ given by Theorem 4.3.
Proposition 5.5.
The category of parahoric $\mathcal{G}_{\boldsymbol{\alpha}}$-torsors on $X$ and the category of $G$-bundles on $\mathscr{X}$ of type $\boldsymbol{\rho}$ are equivalent.
Proof.
We prove the correspondence by discussing transition functions. We first recall some notations. Let $R_{i}$ be the discrete valuation rings given by the generic point $\eta_{i}$ of $D_{i}$. We have
$\mathscr{X}\times_{X}{\rm Spec}\,R_{i}\cong[{\rm Spec}\,\widetilde{R}_{i}/\mu_{d_{i}}].$
By Lemma 4.6, a $G$-bundle $\mathscr{E}$ of type $\boldsymbol{\rho}$ on $\mathscr{X}$ is equivalent to a compatible tuple $(\mathscr{E}_{0},\widehat{\mathscr{E}}_{i},\Theta_{i})$, where
$\Theta_{i}:\widehat{\mathscr{E}}_{i}\times_{[{\rm Spec}\,\widehat{\widetilde{R}}_{i}/\mu_{d_{i}}]}[{\rm Spec}\,\widehat{\widetilde{K}}_{i}/\mu_{d_{i}}]\xrightarrow{\cong}\mathscr{E}_{0}\times_{\mathscr{X}\backslash\widetilde{D}}[{\rm Spec}\,\widehat{\widetilde{K}}_{i}/\mu_{d_{i}}]$
is an isomorphism. Recall that
$\Theta_{i}(\gamma_{i}\cdot w_{i})=\rho_{i}(\gamma_{i})\Theta_{i}(w_{i}).$
We can choose an element $\Delta_{i}\in T(\widehat{\widetilde{R}}_{i})$ such that
$\Delta_{i}(\gamma_{i}\cdot w_{i})=\rho_{i}(\gamma_{i})\Delta_{i}(w_{i}),$
which can be regarded as $\Delta_{i}=w_{i}^{d_{i}\alpha_{i}}$ (see [BS15, (2.2.9.4)]), and define new isomorphisms
$\Theta^{\prime}_{i}:=\Delta^{-1}_{i}\Theta_{i},\quad 1\leq i\leq s,$
and it is easy to check that $\Theta^{\prime}_{i}$ is $\mu_{d_{i}}$-invariant, i.e.
$\Theta^{\prime}_{i}(\gamma_{i}\cdot w_{i})=\Theta^{\prime}_{i}(w_{i}).$
Thus, $\Theta^{\prime}_{i}$ descents to a well-defined element in $G(\widehat{K}_{i})$ under the substitution $z_{i}=w_{i}^{d_{i}}$. Thus, there is a one-to-one correspondence between elements in $G(\widehat{\widetilde{K}}_{i})^{\rho_{i}}$ and elements in $G(\widehat{K}_{i})$.
With a similar approach as [BS15, Theorem 2.3.1], we have
$G(\widehat{\widetilde{R}}_{i})^{\rho_{d_{i}}}\cong G_{\alpha_{i}}(\widehat{R}_{i}).$
More precisely, given an isomorphism $\Psi:\mathscr{E}_{i}\rightarrow\mathscr{E}_{i}$, we regard it as an element in $G(\widehat{\widetilde{R}}_{i})^{\rho_{i}}$, i.e.
$\rho(\gamma_{i})\Psi_{i}(w_{i})\rho(\gamma_{i})^{-1}=\Psi_{i}(\gamma_{i}\cdot w_{i}).$
Define $\Psi^{\prime}_{i}=\Delta_{i}^{-1}\Psi_{i}\Delta_{i}$, which is $\mu_{d_{i}}$-invariant, i.e.
$\Psi^{\prime}_{i}(\gamma_{i}\cdot w_{i})=\Psi^{\prime}_{i}(w_{i}).$
Thus, $\Psi^{\prime}_{i}(w_{i})$ descents to a well-defined element in $G(\widehat{K}_{i})$ and it is easy to check that it is in $G_{\alpha_{i}}(\widehat{R}_{i})$. Indeed, conjugating by $\Delta_{i}$ and substituting $z_{i}=w_{i}^{d_{i}}$ induces the desired isomorphism.
It is easy to check that the compatbility condition for $G$-bundles on root stacks is equivalent to that for parahoric $\mathcal{G}_{\boldsymbol{\alpha}}$-torsors. This finishes the proof of this proposition. ∎
Now we suppose that $X$ is projective. Let $P$ be a parabolic subgroup of $G$. We can define a subgroup scheme $\mathcal{P}_{\boldsymbol{\alpha}}\subseteq\mathcal{G}_{\boldsymbol{\alpha}}$ with respect to a given collection of weights $\boldsymbol{\alpha}$ (see [KSZ24, §4.1]). Given a parahoric $\mathcal{G}_{\boldsymbol{\alpha}}$-torsor $\mathcal{E}$, a reduction of structure group is a section $\varsigma:X\rightarrow\mathcal{E}/\mathcal{P}_{\boldsymbol{\alpha}}$, and then we obtain a parahoric $\mathcal{P}_{\boldsymbol{\alpha}}$-torsor $\mathcal{E}_{\varsigma}$.
${\mathcal{E}_{\varsigma}}$${\mathcal{E}}$${X}$${\mathcal{E}/\mathcal{P}_{\boldsymbol{\alpha}}}$$\scriptstyle{\varsigma}$
Let $\kappa:\mathcal{P}_{\boldsymbol{\alpha}}\rightarrow\mathbb{G}_{m}$ be a morphism of group schemes over $X$, where we use the subscript $X$ to emphasize that $\mathbb{G}_{m}$ is regarded as a group scheme over $X$. This morphism $\kappa$ called a character. Then, we obtain a line bundle $L(\varsigma,\kappa):=\kappa_{*}\mathcal{E}_{\varsigma}$. Moreover, since a character $\kappa:\mathcal{P}_{\boldsymbol{\alpha}}\rightarrow\mathbb{G}_{m}$ can be regarded as a character $\chi:P\rightarrow\mathbb{G}_{m}$ [KSZ24, Lemma 4.1], the pairing $\langle\boldsymbol{\alpha},\kappa\rangle$ is defined as
$\langle\boldsymbol{\alpha},\kappa\rangle:=\langle\boldsymbol{\alpha},\chi\rangle=\sum_{i=1}^{s}\langle\alpha_{i},\chi\rangle.$
Definition 5.6 (Definition 4.2 in [KSZ24]).
Let $\mathcal{E}$ be a parahoric $\mathcal{G}_{\boldsymbol{\alpha}}$-torsor. Given a reduction of structure group $\varsigma$ and a character $\kappa$, we define the parahoric degree as
$par\deg\mathcal{E}(\varsigma,\kappa)=\deg L(\varsigma,\kappa)+\langle\boldsymbol{\alpha},\kappa\rangle.$
We remind the reader that we fix an ample line bundle $H$ on $X$ and the degree of $\deg L(\varsigma,\kappa)$ is given with respect to $H$. Given the definition of parahoric degree, we define the $R$-stability condition for parahoric torsors.
Definition 5.7.
A parahoric $\mathcal{G}_{\boldsymbol{\alpha}}$-torsor $\mathcal{E}$ is $R$-semistable (resp. $R$-stable), if
  • for any proper parabolic subgroup $P\subseteq G$;
  • for any antidominant character $\kappa:\mathcal{P}_{\boldsymbol{\alpha}}\rightarrow\mathbb{G}_{m}$ acting trivially on the center;
  • for any reduction of structure group $\varsigma:X\rightarrow\mathcal{E}/\mathcal{P}_{\boldsymbol{\alpha}}$,
we have
[…]
Let $\mathcal{E}$ be a parahoric $\mathcal{G}_{\boldsymbol{\alpha}}$-torsor on $X$, and denote by $\mathscr{E}$ the corresponding $G$-bundle on $\mathscr{X}$ by Proposition 5.5. The results in [KSZ24] implies the following one-to-one correspondences:
  1. (1)
    characters: $\chi:P\rightarrow\mathbb{G}_{m}$ and $\kappa:P_{\boldsymbol{\alpha}}\rightarrow\mathbb{G}_{m}$ [KSZ24, Lemma 4.1],
  2. (2)
    reductions: $\varsigma:X\rightarrow\mathcal{E}/\mathcal{P}_{\boldsymbol{\alpha}}$ and reductions of $\sigma:\mathscr{X}\rightarrow\mathscr{E}/P$ [KSZ24, Lemma 4.10].
Lemma 5.8.
Under the correspondence of characters and reductions of structure group above, there exists a locally free sheaf $\mathscr{L}$ on $\mathscr{X}$ such that
$\deg_{\mathscr{L}}L(\sigma,\chi)=(d_{1}\dots d_{s})par\deg\mathcal{E}(\varsigma,\kappa).$
Proof.
[…]
$\mathscr{L}:=\sum_{i=1}^{d_{1}}\mathcal{O}_{\mathscr{X}}(i\widetilde{D}_{1}).$
Borne [Bor07, Theorem 4.3] proved the formula
$\deg_{\mathscr{L}}L(\sigma,\chi)=d_{1}(\deg L(\varsigma,\kappa)+\langle\boldsymbol{\alpha},\kappa\rangle).$
Thus,
$\deg_{\mathscr{L}}L(\sigma,\chi)=d_{1}par\deg\mathcal{E}(\varsigma,\kappa).$
This finishes the proof of this lemma. ∎
Proposition 5.9.
Let $\mathcal{E}$ be a parahoric $\mathcal{G}_{\boldsymbol{\alpha}}$-torsor on $X$. Denote by $\mathscr{E}$ the corresponding $G$-bundle of type $\boldsymbol{\rho}$ on $\mathscr{X}$. Then $\mathcal{E}$ is $R$-(semi)stable if and only if $\mathscr{E}$ is $R$-(semi)stable. Thus, the category of $R$-(semi)stable parahoric $\mathcal{G}_{\boldsymbol{\alpha}}$-torsors on $X$ is equivalent to the category of $R$-(semi)stable $G$-bundle of type $\boldsymbol{\rho}$ on $\mathscr{X}$.
Proof.
The proof of this theorem is the same as that of [KSZ24, Theorem 4.12]. Since the characters $\chi:P\rightarrow\mathbb{G}_{m}$ and reductions of structure group $\sigma:\mathscr{X}\rightarrow\mathscr{E}/P$ are in one-to-one correspondence with characters $\kappa:\mathcal{P}_{\boldsymbol{\alpha}}\rightarrow\mathbb{G}_{m}$ and reductions $\varsigma:X\rightarrow\mathcal{E}/\mathcal{P}_{\boldsymbol{\alpha}}$ respectively, the theorem follows directly from the equality of the corresponding degree by Lemma 5.8. ∎
5.3. Logahoric Higgs Torsors and Logahoric Connections
Definition 5.10.
A logahoric $\mathcal{G}_{\boldsymbol{\alpha}}$-Higgs torsor is a pair $(\mathcal{E},\vartheta)$, where $\mathcal{E}$ is a parahoric $\mathcal{G}_{\boldsymbol{\alpha}}$-torsor and $\vartheta:X\rightarrow\mathcal{E}(\mathfrak{g})\otimes\Omega_{X}({\rm log}\,D)$ is a logarithmic Higgs field. Moreover, $(\mathcal{E},\vartheta)$ is $R$-semistable (resp. $R$-stable) if
  • for any proper parabolic subgroup $P\subseteq G$;
  • for any antidominant character $\kappa:\mathcal{P}_{\boldsymbol{\alpha}}\rightarrow\mathbb{G}_{m}$ acting trivially on the center;
  • for any $\vartheta$-compatible reduction of structure group $\varsigma:X\rightarrow\mathcal{E}/\mathcal{P}_{\boldsymbol{\alpha}}$,
we have
[…]
Proposition 5.11.
Logahoric $\mathcal{G}_{\boldsymbol{\alpha}}$-Higgs torsors $(\mathcal{E},\vartheta)$ on $X$ are in one-to-one correspondence with logarithmic $G$-Higgs bundles $(\mathscr{E},\theta)$ of type $\boldsymbol{\rho}$ on $\mathscr{X}$. Moreover, this correspondence preserves the stability conditions. Thus, the category of $R$-(semi)stable logahoric $\mathcal{G}_{\boldsymbol{\alpha}}$-Higgs torsors on $X$ is equivalent to the category of $R$-(semi)stable logarithmic $G$-Higgs bundles of type $\boldsymbol{\rho}$ on $\mathscr{X}$.
Proof.
This proposition is proved in the same way as [KSZ24, Theorem 4.13]. By Proposition 5.5, parahoric $\mathcal{G}_{\boldsymbol{\alpha}}$-torsors are in one-to-one correspondence with $G$-bundles on $\mathscr{X}$ of type $\boldsymbol{\rho}$, and we have to prove such a correspondence also holds for Higgs fields. It is enough to work on local charts $[{\rm Spec}\,\widetilde{R}_{i}/\mu_{d_{i}}]$ and ${\rm Spec}\,R_{i}$, and we follow the same notation as in §4.2. The restriction of the Higgs field $\vartheta_{i}$ on $\mathscr{E}_{i}$ is $\mu_{d_{i}}$-equivariant, i.e.
$\vartheta_{i}(\gamma_{i}w_{i})=\rho(\gamma_{i})\vartheta_{i}(w_{i})\rho(\gamma_{i})^{-1}.$
Define $\theta_{i}:=\Delta_{i}^{-1}\vartheta_{i}\Delta_{i}$, and the new element $\theta_{i}$ is $\mu_{d_{i}}$-invariant, i.e.
$\theta_{i}(\gamma_{i}w_{i})=\theta_{i}(w_{i}).$
Thus, $\theta_{i}(z_{i})$ is well-defined by taking $z_{i}=w_{i}^{d_{i}}$. With the same argument as in Proposition 5.5, the new element $\theta_{i}(z_{i})$ gives a section ${\rm Spec}\,R_{i}\rightarrow\mathcal{E}(\mathfrak{g})\otimes\Omega_{X}({\rm log}D)|_{{\rm Spec}\,R_{i}}$. Then, Higgs fields $\vartheta$ of $\mathcal{E}$ are in one-to-one correspondence with Higgs fields $\theta$ of $\mathscr{E}$. ∎
We define nilpotent logahoric Higgs torsors in the same way as that for $G$-Higgs bundles in §2.2, and we introduce the following categories:
  • ${\rm HIG}_{p-1}(X_{\rm log}/k,\mathcal{G}_{\boldsymbol{\alpha}})$: the category of nilpotent logahoric $\mathcal{G}_{\boldsymbol{\alpha}}$-Higgs torsors on $X$ of exponent $\leq p-1$;
  • ${\rm HIG}_{p-1}(\mathscr{X}_{\rm log}/k,G,\boldsymbol{\rho})$: the category of nilpotent logarithmic $G$-Higgs bundles of type $\boldsymbol{\rho}$ on $\mathscr{X}$ of exponent $\leq p-1$;
As a direct result of Proposition 5.11, we have the desired equivalence of categories:
Corollary 5.12.
The following categories are equivalent:
  • ${\rm HIG}_{p-1}(\mathscr{X}_{\rm log}/k,G,\boldsymbol{\rho})$ and ${\rm HIG}_{p-1}(X_{\rm log}/k,\mathcal{G}_{\boldsymbol{\alpha}})$;
  • ${\rm HIG}^{(s)s}_{p-1}(\mathscr{X}_{\rm log}/k,G,\boldsymbol{\rho})$ and ${\rm HIG}^{(s)s}_{p-1}(X_{\rm log}/k,\mathcal{G}_{\boldsymbol{\alpha}})$.
Now we consider logarithmic connections. Chen–Zhu gave the definition of (integrable) connections on $\mathcal{G}$-torsors, where $\mathcal{G}$ is an arbitrary smooth affine group scheme over $X$ [CZ15, Appendix]. We follow their setup and define logarithmic integrable connections on parahoric torsors. Let $\mathcal{V}$ be a parahoric $\mathcal{G}_{\boldsymbol{\alpha}}$-torsor. A (integrable) logarithmic connection $\nabla^{\prime}$ on $\mathcal{V}$ is a (integrable) connection $\nabla^{\prime}:\mathcal{O}_{\mathcal{V}}\rightarrow\mathcal{O}_{\mathcal{V}}\otimes_{\mathcal{O}_{X}}\Omega_{X}({\rm log}\,D)$ such that the following diagram commutes
${\mathcal{O}_{\mathcal{V}}}$${\mathcal{O}_{\mathcal{V}}\otimes_{\mathcal{O}_{X}}\mathcal{O}_{G\times X}}$${\mathcal{O}_{\mathcal{V}}\otimes_{\mathcal{O}_{X}}\Omega_{X}({\rm log}\,D)}$${(\mathcal{O}_{\mathcal{V}}\otimes_{\mathcal{O}_{X}}\mathcal{O}_{G\times X})\otimes_{\mathcal{O}_{X}}\Omega_{X}({\rm log}\,D)}$$\scriptstyle{\nabla^{\prime}}$$\scriptstyle{\nabla^{\prime}\otimes 1+1\otimes\nabla_{\mathcal{G}_{\boldsymbol{\alpha}}}}$
where $\nabla_{\mathcal{G}_{\boldsymbol{\alpha}}}$ is an integrable connection on $\mathcal{O}_{\mathcal{G}_{\boldsymbol{\alpha}}}$.
Definition 5.13.
A logahoric $\mathcal{G}_{\boldsymbol{\alpha}}$-connection is a pair $(\mathcal{V},\nabla^{\prime})$, where $\mathcal{V}$ is a parahoric $\mathcal{G}_{\boldsymbol{\alpha}}$-torsor and $\nabla^{\prime}:\mathcal{O}_{\mathcal{V}}\rightarrow\mathcal{O}_{\mathcal{V}}\otimes\Omega_{X}({\rm log}\,D)$ is a logarithmic integrable connection.
Given a logahoric $\mathcal{G}_{\boldsymbol{\alpha}}$-connection, denote by $\psi^{\prime}:=\psi(\nabla^{\prime})$ the $p$-curvature of $\nabla^{\prime}$. Then, the nilpotency of a logarithmic integrable connection $\nabla^{\prime}$ is defined in the same way as §2.2. Furthermore, we define the residue map
${\rm Res}_{D_{i}}\nabla^{\prime}:\mathcal{O}_{\mathcal{V}}\otimes\mathcal{O}_{D_{i}}\rightarrow\mathcal{O}_{\mathcal{V}}\otimes\mathcal{O}_{D_{i}}$
as in §4.3. We introduce the following categories:
  • ${\rm MIC}_{p-1}(X_{\rm log}/k,\mathcal{G}_{\boldsymbol{\alpha}})$: the category of nilpotent logahoric $\mathcal{G}_{\boldsymbol{\alpha}}$-connections $(\mathcal{V},\nabla^{\prime})$ of exponent $\leq p-1$ on $X$ such that
    1. (1)
      the semisimple part of the residue ${\rm Res}_{D_{i}}\nabla^{\prime}$ is $\alpha_{i}$ for each $1\leq i\leq s$, where we regard $\alpha_{i}$ as an element in the Lie algebra $\mathfrak{g}$;
    2. (2)
      the nilpotent part of ${\rm Res}_{D_{i}}\nabla^{\prime}$ is nilpotent of exponent $\leq p-1$ for every $1\leq i\leq s$.
  • ${\rm MIC}_{p-1}(\mathscr{X}_{\rm log}/k,G,\boldsymbol{\rho})$: the category of nilpotent flat $G$-bundles $(\mathscr{V},\nabla)$ of type $\boldsymbol{\rho}$ of exponent $\leq p-1$ on $\mathscr{X}$, of which the residue ${\rm Res}_{\widetilde{D}_{i}}\nabla$ is nilpotent of exponent $\leq p-1$ for every $1\leq i\leq s$.
The extra condition that the semisimple part of the residue ${\rm Res}_{D_{i}}\nabla^{\prime}$ is $\alpha_{i}$ for each $1\leq i\leq s$ in the definition of the category ${\rm MIC}_{p-1}(X_{\rm log}/k,\mathcal{G}_{\boldsymbol{\alpha}})$ is called adjustness, which is studied by the authors in [KS20, §2].
Proposition 5.14.
The categories ${\rm MIC}_{p-1}(X_{\rm log}/k,\mathcal{G}_{\boldsymbol{\alpha}})$ and ${\rm MIC}_{p-1}(\mathscr{X}_{\rm log}/k,G,\boldsymbol{\rho})$ are equivalent.
Proof.
The discussion is similar to that of Higgs bundles (Proposition 5.11). We still work on local charts $[{\rm Spec}\,\widetilde{R}_{i}/\mu_{d_{i}}]$ and ${\rm Spec}\,R_{i}$. The only thing we have to show is how the residues of connections change. Let $(\mathscr{V},\nabla)$ be a logarithmic flat $G$-bundle on $\mathscr{X}$. Recall that the transition function $\Theta_{i}$ changes to $\Theta^{\prime}_{i}=\Delta_{i}^{-1}\Theta_{i}$ (Proposition 5.5), which results in the change of Higgs fields $\theta_{i}=\Delta_{i}^{-1}\vartheta_{i}\Delta_{i}$ (Proposition 5.11). Therefore, we define a new connection $\nabla^{\prime}_{i}(w_{i}):=\Delta_{i}(w_{i})\circ\nabla_{i}(w_{i})$, where the action is the gauge action. Regarding the element $\Delta_{i}(w_{i})$ as $w_{i}^{d_{i}\alpha_{i}}$, we have
$\nabla^{\prime}_{i}(w_{i})=d_{i}\alpha_{i}\frac{dw_{i}}{w_{i}}+\Delta_{i}^{-1}(w_{i})\nabla_{i}(w_{i})\Delta_{i}(w_{i}),$
which is $\mu_{d_{i}}$-invariant. Under the substitution $z_{i}=w_{i}^{d_{i}}$, we obtain a connection $\nabla^{\prime}_{i}(z_{i})$. Since the residue of $\nabla_{i}$ is nilpotent, the semisimple part of $\nabla^{\prime}_{i}(z_{i})$ is $\alpha_{i}$. ∎
Definition 5.15.
Let $(\mathcal{V},\nabla^{\prime})$ be a logahoric $\mathcal{G}_{\boldsymbol{\alpha}}$-connection. It is $R$-semistable (resp. $R$-stable) if
  • for any proper parabolic subgroup $P\subseteq G$;
  • for any antidominant character $\kappa:\mathcal{P}_{\boldsymbol{\alpha}}\rightarrow\mathbb{G}_{m}$ acting trivially on the center;
  • for any $\nabla^{\prime}$-compatible reduction of structure group $\varsigma:X\rightarrow\mathcal{V}/\mathcal{P}_{\boldsymbol{\alpha}}$,
we have
[…]
Adding the stability conditions, we have:
Corollary 5.16.
The categories ${\rm MIC}^{(s)s}_{p-1}(X_{\rm log}/k,\mathcal{G}_{\boldsymbol{\alpha}})$ and ${\rm MIC}^{(s)s}_{p-1}(\mathscr{X}_{\rm log}/k,G,\boldsymbol{\rho})$ are equivalent.
Denote by $\boldsymbol{\rho}$ the collection of corresponding representations of $\boldsymbol{\alpha}$. Let $\boldsymbol{\rho}^{p}$ be the collection of representations given in Remark 4.13 and denote by $\boldsymbol{\alpha}^{\prime}$ the corresponding collection of weights. In conclusion, we obtain the logahoric version of nonabelian Hodge correspondence:
Theorem 5.17.
Suppose that $(X,D)$ is $W_{2}(k)$-liftable. The following categories are equivalent:
  • ${\rm HIG}_{p-1}(X^{\prime}_{\rm log}/k,\mathcal{G}_{\boldsymbol{\alpha}})$ and ${\rm MIC}_{p-1}(X_{\rm log}/k,\mathcal{G}_{\boldsymbol{\alpha}^{\prime}})$,
  • ${\rm HIG}^{(s)s}_{p-1}(X^{\prime}_{\rm log}/k,\mathcal{G}_{\boldsymbol{\alpha}})$ and ${\rm MIC}^{(s)s}_{p-1}(X_{\rm log}/k,\mathcal{G}_{\boldsymbol{\alpha}^{\prime}})$.
Proof.
The proof of the first equivalence follows directly from the following diagram,
[…]
and the second equivalence is given via the following diagram.
[…]
The only thing we have to prove is that if $(X,D)$ is $W_{2}(k)$-liftable, the corresponding pairs $(\mathscr{X},\widetilde{D})$ is also $W_{2}(k)$-liftable, where $\mathscr{X}$ is the root stack determined by the data $(X,D,\boldsymbol{d})$. By the well-known covering lemma of Kawamata [Kaw81, Theorem 17], there exists a covering $\pi^{\prime}:Y\rightarrow X$ such that $Y$ is smooth and $D^{\prime}:=(\pi^{\prime*}D)_{red}$ is a reduced normal crossing divisor on $Y$. Moreover, we can assume that the induced morphism $Y\rightarrow\mathscr{X}$ is surjective and étale [Sim11, Lemma 4.2]. Note that although Kawamata and Simpson stated the results only in characteristic zero, their arguments hold in our case due to the assumption that all integers in $\boldsymbol{d}$ are coprime to $p$. Then it is enough to show that the pair $(Y,D^{\prime})$ is $W_{2}(k)$-liftable. This property has been proved in [XW13, Theorem 1.1] and we also refer the reader to [KS20, Notation 2.16] for the case of curves. ∎
Remark 5.18.
When $X$ is a curve and $G={\rm GL}_{n}$, the correspondence given in Theorem 5.17 has been studied in [KS20, Theorem 2.10], where the author used parabolic Higgs bundles and parabolic flat bundles to prove this result.
Acknowledgement. The first and third named authors are partially supported by the National Key R and D Program of China 2020YFA0713100, CAS Project for Young Scientists in Basic Research Grant No. YSBR-032. The second named author is supported by Guangdong Basic and Applied Basic Research Foundation 2024A04J0334 and 2024A1515011583.